### v4

# The Identification Power of

Abstract
This study investigates the identification power gained by combining experimental data, in which treatment is randomized, with observational data, in which treatment is self-selected, for distributional treatment effect (DTE) parameters. While experimental data identify average treatment effects, many DTE parameters—such as the distribution of individual treatment effects—are only partially identified. We examine whether and how combining these two data sources tightens the identified set for such parameters. For broad classes of DTE parameters, we derive nonparametric sharp bounds under the combined data and clarify the mechanism through which data combination improves identification relative to using experimental data alone. Our analysis highlights that self-selection in observational data is a key source of identification power. We establish necessary and sufficient conditions under which the combined data shrink the identified set, showing that such shrinkage generally occurs unless selection-on-observables holds in the observational data. We also propose a linear programming approach to compute sharp bounds that can incorporate additional structural restrictions, such as positive dependence between potential outcomes and the generalized Roy selection model. An empirical application using data on negative campaign advertisements in the 2008 U.S. presidential election illustrates the practical relevance of the proposed approach.
Keywords: Data combination; Distributional treatment effect; Heterogeneous treatment effect; Partial identification; Self-selection.
1 Introduction
Researchers often have access to both experimental and observational data when evaluating public policies and medical interventions. Experimental studies, such as randomized controlled trials, are regarded as the gold standard for causal inference, whereas observational data are far more prevalent in practice and often contain rich behavioral variation. This study investigates the identification gains that result from combining these two data sources, with a focus on distributional treatment effects (DTEs).
We consider a general setting in which treatment receipt and outcomes are observed from two data sources: experimental data, where the treatment is randomly assigned, and observational data, where the treatment is self-selected. This structure arises naturally in various policy and medical contexts. In development policy, for example, the efficacy of public health interventions such as mosquito nets is evaluated through field experiments (where nets are randomly assigned) and observational data (where nets are self-purchased). In education, randomized class-size assignments (e.g., Project STAR) coexist with parental or administrative decisions regarding classroom placement. In clinical medicine, drugs may be randomly assigned in clinical trials; however, once approved, they are prescribed based on a physician’s judgment.
Experimental data are valuable because the random assignment of treatment enables identification of key causal parameters such as the average treatment effect (ATE). However, beyond the ATE, DTE parameters are often crucial to policy and medical evaluation, as well as for understanding treatment effect heterogeneity. Let $Y_{1}$ and $Y_{0}$ denote potential outcomes under treatment and no-treatment, respectively. Examples of such parameters include (i) the fraction of individuals who benefit from treatment $\mathbb{P}(Y_{1}>Y_{0})$; (ii) distribution function of individual treatment effects $F_{Y_{1}-Y_{0}}(\delta)\equiv\mathbb{P}(Y_{1}-Y_{0}\leq\delta)$; (iii) ATE for disadvantaged individuals $\mathbb{E}[Y_{1}-Y_{0}\mid Y_{0}\leq c]$, where the subpopulation with $Y_{0}\leq c$ represents individuals whose baseline outcomes fall below a threshold level $c$ (e.g., a poverty line); and (iv) the correlation between $Y_{1}$ and $Y_{0}$.
Although these parameters are highly relevant for policy and medical evaluations, they are generally not point-identified from experimental data alone because their identification requires knowledge of the joint distribution of the two potential outcomes. Experimental data identify only the marginal distributions of $Y_{1}$ and $Y_{0}$ and hence yield partial identification (e.g., Heckman et al. (1997); Fan and Park (2010)). However, the resulting identified sets are often wide and not very informative.
This study explores whether and how combining experimental and observational data can improve the identification of distributional parameters. Specifically, we address the following questions. (i) Does combining the two data sources shrink the identified set for DTE parameters? (ii) If so, what mechanism drives this shrinkage, and under what conditions does it occur? (iii) How can the identified set under the combined data be characterized and computed?
To address these questions, we build on a framework that incorporates both experimental and observational data. In this setting, researchers observe individuals’ outcomes $Y$, treatment status $D\in\{0,1\}$, and a data source indicator $G\in\{\mathrm{exp},\mathrm{obs}\}$, which indicates whether an observation comes from an experimental study ($G=\mathrm{exp}$) or an observational study ($G=\mathrm{obs}$). A key element of the framework is the latent self-selection type $S\in\{0,1\}$, which represents the treatment choice an individual would make under self-selection. This variable is observed in the observational data (where $S=D$) but unobserved in the experimental data.
We define the identified sets under experimental data alone and under the combined data, and then use a copula-based approach to characterize and compare them. Under standard assumptions—random treatment assignment in the experimental data and external validity across data sources—this framework enables a systematic comparison of what can be learned from experimental data alone versus combining both data sources.
Our central theoretical results show that the identified set under the combined data is generally smaller than that under the experimental data alone, except in special cases where the latent self-selection $S$ is independent of the potential outcomes. The intuition is as follows. While the experimental data identify only the marginal distributions, $F_{Y_{1}}$ and $F_{Y_{0}}$, of the potential outcomes, combining the two data sources enables identification of the joint distributions $(F_{Y_{1}S},F_{Y_{0}S})$ of each potential outcome and the self-selection type. We show that the distribution pairs $(F_{Y_{1}},F_{Y_{0}})$ and $(F_{Y_{1}S},F_{Y_{0}S})$ fully characterize the identified sets under the experimental and combined data, respectively. As the latent self-selection $S$ encodes information about the dependence structure between the potential outcomes, knowledge of $(F_{Y_{1}S},F_{Y_{0}S})$ can tighten the set of feasible joint distributions of $(Y_{1},Y_{0})$.
To quantify this identification gain and derive nonparametric sharp bounds, we apply copula bound analysis that builds on and extends the work of Fan et al. (2017), who study the identifying power of covariates, to settings with the self-selection variable $S$. For broad classes of DTE parameters represented by supermodular functions or $\varphi$-indicator functions (defined later), we derive analytical expressions for the sharp bounds and establish necessary and sufficient conditions under which the data combination yields tighter identified sets. We show that combining the two data sources improves identification when the dependence structure between the potential outcomes varies across the latent self-selection types.
To further tighten the identified sets, we also consider incorporating additional structural restrictions that are plausible in many empirical contexts. In particular, we consider two such restrictions: positive dependence between the potential outcomes (Joe, 2014; Frandsen and Lefgren, 2021) and a generalized Roy selection model. Though not required for our main results, these restrictions can substantially narrow the identified sets. To implement these restrictions under the data combination, we develop a linear programming approach that efficiently computes sharp bounds for a broad class of DTE parameters.
Our analysis also extends to data from doubly randomized preference trials (DRPTs) (Rücker, 1989; Long et al., 2008), a unique yet recently prevalent design. In DRPTs, individuals are randomly assigned to one of three groups: a treatment group, a control (no-treatment) group, or a self-selection group, in which individuals choose between treatment and no-treatment. This design has been used in a wide range of studies across the social sciences (Gaines and Kuklinski, 2011; Arceneaux et al., 2012; De Benedictis-Kessner et al., 2019; Ida et al., 2025) and medical sciences (King et al., 2005; Howard and Thornicroft, 2006). An established advantage of DRPTs is that they enable identification of the ATE for each self-selection group, $\mathbb{E}[Y_{1}-Y_{0}\mid S=s]$ for $s\in\{0,1\}$ (Long et al., 2008). Our results highlight a novel advantage of this design: it improves the identification of DTEs by combining random-assignment and self-selection samples.
We illustrate the empirical relevance of our approach using DRPT data from Gaines and Kuklinski (2011), who study the effects of negative campaign advertisements on individuals’ attitudes toward presidential candidates during the 2008 U.S. presidential election. We find that incorporating self-selection data substantially tightens the identified sets for DTE parameters such as $\mathbb{P}(Y_{1}<Y_{0})$, the fraction of individuals negatively affected by the advertisements. The results further indicate that the advertisements have substantial but heterogeneous effects. These findings highlight the empirical value of combining random-assignment and self-selection data for estimating DTEs.
Related Literature
This study relates and contributes to two strands of literature: (i) the combination of random-assignment and self-selection data sources for causal inference, and (ii) partial or point identification of DTEs.
The first strand explores the benefits of combining data from random-assignment and self-selection sources. This literature typically pursues two objectives: (1) improving the precision of estimates when experimental data suffice for identification and (2) identifying causal effects that are not identifiable from experimental data alone. Regarding the first objective, Rosenman et al. (2023), Yang et al. (2023), and Gui (2024) propose novel estimation methods that improve the efficiency of treatment effect estimation by combining experimental and observational data. Regarding the second objective, Long et al. (2008) show that the ATE for each self-selection group, $\mathbb{E}[Y_{1}-Y_{0}|S=s]$ ($s=0,1$), can be identified using data from a DRPT. Knox et al. (2019) extend this to multiple treatment settings and derive partial identification results. In a different context, where experimental data contain secondary outcomes and observational data contain primary outcomes, Athey et al. (2025) study identification of the ATE for the primary outcomes. Our setting differs from theirs in that we consider two data sources sharing the same outcome type.
Our contribution to this strand of the literature is to uncover a novel advantage of data combination: it can shrink the identified sets for DTE parameters. Unlike prior studies that focus on point identification of new parameters or gains in estimation efficiency, we show that combining data improves informativeness by tightening bounds in partially identified settings.
The second strand studies the partial or point identification of DTEs under various assumptions and structural restrictions. [^1] Notable works include Heckman et al. (1997), Manski (1997), Fan and Park (2009; 2010; 2012), Fan and Wu (2010), Fan et al. (2014), Fan et al. (2017), Vuong and Xu (2017), Kim et al. (2018), Firpo and Ridder (2019), Mourifie et al. (2020), Callaway (2021), Frandsen and Lefgren (2021), Russell (2021), Lee (2024), Cui and Han (2025), and Kaji and Cao (2025) among others. In particular, Fan and Park (2009; 2010; 2012), Fan et al. (2017), and Firpo and Ridder (2019) develop nonparametric bounds for the joint distribution of potential outcomes and its functionals under minimal assumptions such as random treatment assignment. To tighten these bounds, subsequent studies have explored additional structures, including self-selection models (Fan and Wu, 2010; Mourifie et al., 2020), panel data with time-dependence restrictions (Callaway, 2021), mutual stochastic monotonicity of potential outcomes (Frandsen and Lefgren, 2021), and so on.
This study contributes to this literature by clarifying the identifying power that arises from combining experimental and observational data—a novel structure not previously explored—and by providing computable characterizations of the sharp bounds under the data combination.
Finally, this study contributes to the broader causal inference literature by offering a new perspective on the role of observational data in identification. While the prevailing view holds that observational data become redundant for identification once experimental data are available, this study shows that they can nonetheless enhance the identification of DTEs, even in the presence of experimental data.
Structure of the Paper
The remainder of the paper is organized as follows. Section 2 introduces the data combination framework, formalizes the DTE parameters, and defines their identified sets under experimental data alone and under the combined data. Section 3 characterizes the identified sets under each data scenario and highlights the sources of identification gains from data combination. Section 4 derives sharp bounds for broad classes of DTE parameters, specifically those represented by super-modular or $\varphi$-indicator functions, and establishes necessary and sufficient conditions under which data combination improves identification. Section 5 introduces additional restrictions and develops a linear programming approach to compute sharp bounds. Section 6 provides an empirical illustration using DRPT data from Gaines and Kuklinski (2011). Section 7 concludes. All proofs and numerical examples are provided in the appendix.
2 Setup
We begin by outlining the framework. Section 2.1 introduces the data combination setting and fundamental assumptions. Section 2.2 defines the DTE parameters and provides illustrative examples. Section 2.3 defines the identified sets under experimental and combined data.
2.1 Combined Experimental and Observational Data
We consider observed data consisting of the quadruple $(Y,D,G,X)$, where $Y$ is the outcome; $D\in\{0,1\}$ is a binary treatment indicator; $G\in\{\mathrm{exp},\mathrm{obs}\}$ denotes the data source; and $X$ is a vector of pre-treatment covariates with its support denoted by $\mathcal{X}$. Specifically, $G$ indicates whether an observation comes from an experimental study ($G=\mathrm{exp}$) or observational study ($G=\mathrm{obs}$). The treatment indicator $D$ equals $1$ if the individual receives treatment and 0 otherwise. In the experimental data ($G=\mathrm{exp}$), $D$ is randomly assigned, whereas in the observational data ($G=\mathrm{obs}$), it is determined through self-selection. Let $Y_{1}$ and $Y_{0}$ denote the potential outcomes under treatment and no-treatment, respectively, both of which are assumed to be continuous. The observed outcome is defined as $Y\equiv DY_{1}+(1-D)Y_{0}$.
We introduce a latent self-selection variable $S\in\{0,1\}$ that represents the treatment choice an individual would make if he or she self-selects. In the observational data ($G=\mathrm{obs}$), $S$ coincides with the observed treatment, i.e., $S=D$, whereas in the experimental data ($G=\mathrm{exp}$), $S$ is unobserved. The variable $S$ can also be interpreted as a latent preference for treatment versus no-treatment.
Let $F^{\ast}$ denote the true cumulative distribution function (CDF) of all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$, including both observed and unobserved ones. For a generic CDF $F$, we denote the probability and expectation under $F$ by $\mathbb{P}_{F}(\cdot)$ and $\mathbb{E}_{F}[\cdot]$, respectively. For the true CDF $F^{\ast}$, we use the shorthand $\mathbb{P}(\cdot)$ and $\mathbb{E}[\cdot]$ in place of $\mathbb{P}_{F^{\ast}}(\cdot)$ and $\mathbb{E}_{F^{\ast}}[\cdot]$.
Throughout the paper, we suppose that $F^{\ast}$ satisfies the following assumptions.
Assumption 2.1 (Oberved Outcomes).
$Y=DY_{1}+(1-D)Y_{0}$ a.s.
Assumption 2.2 (Self-selection).
$\mathbb{P}(S=D\mid G=\mathrm{obs},X)=1$ a.s.
Assumption 2.3 (Random Assignment).
$(Y_{1},Y_{0})\perp\!\!\!\perp D\mid X,G=\mathrm{exp}$.
Assumption 2.4 (External Validity).
$(Y_{1},Y_{0},S)\perp\!\!\!\perp G|X$.
Assumption 2.5 (Overlap).
$\mathbb{P}(D=d,G=g|X)>0$ a.s. for all $d\in\{0,1\}$ and $g\in\{\mathrm{exp},\mathrm{obs}\}$.
Assumption 2.2 states that self-selection corresponds to the received treatment in the observational data, which is trivially satisfied by the definition of $S$. Assumption 2.3 requires that treatment is randomly assigned in the experimental study, possibly conditional on $X$. Assumption 2.4 concerns the external validity of the data sources, requiring that any systematic difference between the populations in the experimental and observational studies are captured by $X$. This assumption is automatically satisfied under a DRPT design, where the data source indicator $G$ is randomly assigned. In such a design, individuals are randomly assigned to treatment, no-treatment, and self-selection groups. Assumption 2.5 imposes an overlap condition on both the treatment status and data source. [^2] This overlap condition can be relaxed at the cost of yielding wider identified sets under the combined data. For example, if a covariate value $x$ does not satisfy $\mathbb{P}(D=d,G=g\mid X=x)>0$ for some $g$, then experimental and observational data cannot be combined at that value of $x$. In this case, one may still rely on either the experimental or the observational data alone to construct an identified set conditional on $x$.
2.2 Distributional Treatment Effect Parameters
Following Fan et al. (2017), we consider a parameter of interest $\theta_{o}\in\mathbb{R}$ that has the following form:
$\theta_{o}\equiv\mathbb{E}\left[\psi(Y_{1},Y_{0})\right],$ (1)
where $\psi:\mathbb{R}^{2}\to\mathbb{R}$ is a given function. With various specifications of $\psi$, this formulation encompasses a wide range of DTE parameters, as illustrated below.
Example 2.1 (Fraction of Positive Treatment Effects).
When $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}>y_{0}\}$, we have $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, the fraction of individuals who benefit from the treatment. This parameter captures the distributional impact of the treatment. Importantly, even if the ATE $\mathbb{E}[Y_{1}-Y_{0}]$ is positive, the fraction $\mathbb{P}(Y_{1}>Y_{0})$ may be small, indicating that a small number of individuals experience large gains while many experience no gains or losses. Focusing on $\mathbb{P}(Y_{1}>Y_{0})$ is therefore essential to avoid interventions that are unfavorable for the majority.
Example 2.2 (Distribution Function of Treatment Effects).
More generally, when $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}-y_{0}<\delta\}$ for a fixed $\delta$, the parameter $\theta_{o}$ corresponds to the distribution function of the individual treatment effect: $F_{Y_{1}-Y_{0}}^{\ast}(\delta)\equiv\mathbb{P}(Y_{1}-Y_{0}\leq\delta)$.
Example 2.3 (ATE for the Disadvantaged).
Consider the ATE for disadvantaged individuals, defined as $\mathbb{E}[Y_{1}-Y_{0}\mid Y_{0}\leq c]$ for a fixed threshold $c$. This parameter can be expressed in the form of equation (1) with $\psi(y_{1},y_{0})=(y_{1}-y_{0})\cdot\mathbf{1}\{y_{0}\leq c\}/\mathbb{P}(Y_{0}\leq c)$. The subpopulation with $Y_{0}\leq c$ represents individuals whose outcomes fall below the disadvantage threshold $c$ (e.g., a poverty line) without treatment. The nuisance parameter $\mathbb{P}(Y_{0}\leq c)$ in $\psi(y_{1},y_{0})$ is point-identified from the experimental data under the maintained assumptions as $\mathbb{P}(Y_{0}\leq c)=\mathbb{P}(Y\leq c\mid D=0,G=\textit{exp})$.
Example 2.4 (Probability of Upward Mobility).
When $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}>c,y_{0}\leq c\}/\mathbb{P}(Y_{0}\leq c)$ for a fixed threshold $c$, the target parameter $\theta_{o}$ becomes $\theta_{o}=\mathbb{P}(Y_{1}>c\mid Y_{0}\leq c)$. This parameter represents the probability of upward mobility among individuals whose outcomes would fall below $c$ in the absence of treatment. Specifically, it captures the likelihood that individuals whose baseline outcomes are below a given threshold—such as a poverty line—will exceed that threshold when treated. Unlike the ATE, this parameter focuses on distributional gains among the disadvantaged subpopulation.
Example 2.5 (Correlation between Two Potential Outcomes).
When $\psi(y_{1},y_{0})=(y_{1}-\mathbb{E}[Y_{1}])(y_{0}-\mathbb{E}[Y_{0}])/\sqrt{\text{Var}(Y_{1})\text{Var}(Y_{0})}$, we have $\theta_{o}=\text{Cor}(Y_{1},Y_{0})$, the correlation between the two potential outcomes. This parameter reflects the degree of similarity in individual-level responses to the two interventions, offering insights into treatment effect heterogeneity. The nuisance components $\mathbb{E}[Y_{d}]$ and $\text{Var}(Y_{d})$ for $d\in\{0,1\}$ are point-identified from the experimental data under the maintained assumptions.
See also Fan et al. (2017) for additional examples.
Note that none of the DTE parameters introduced above can be point-identified even with experimental data. Their identification generally requires knowledge of the joint distribution of the two potential outcomes, $F_{Y_{1}Y_{0}}^{\ast}$, whereas experimental data identify only the marginal distributions, $(F_{Y_{1}}^{\ast},F_{Y_{0}}^{\ast})$.
We can also define DTE parameters for various subpopulations: by self-selection type, $\theta_{o,s}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid S=s]$; conditional on covariates, $\theta_{o,x}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid X=x]$; and by data source, $\theta_{o,g}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid G=g]$ for $g\in\{\mathrm{exp},\mathrm{obs}\}$. In particular, $\theta_{o,s}$ captures the treatment effect for individuals who self-select into treatment ($s=1$) or no-treatment ($s=0$), a quantity that is relevant in many empirical applications.
These subpopulation parameters can be incorporated into our framework with slight modifications. For example, the parameter $\theta_{o,g}$ can be expressed as
$\theta_{o,g}=\mathbb{E}\left[\psi(Y_{1},Y_{0})\cdot\frac{\mathbf{1}\{G=g\}}{\mathbb{P}(G=g)}\right],$
where $\mathbb{P}(G=g)$ is point-identified from the observed data. Defining $\tilde{\psi}(y_{1},y_{0})=\psi(y_{1},y_{0})\cdot\mathbf{1}\{G=g\}/\mathbb{P}(G=g)$, we can write $\theta_{o,g}=\mathbb{E}[\tilde{\psi}(Y_{1},Y_{0})]$, so it can be analyzed analogously to $\theta_{o}$. As for $\theta_{o,s}$, Lemma C.1 in the appendix shows that
$\theta_{o,s}=\mathbb{E}[\psi(Y_{1},Y_{0})|D=s,G=\mathrm{obs}]=\mathbb{E}\left[\psi(Y_{1},Y_{0})\cdot\frac{\mathbf{1}\{D=s,G=\mathrm{obs}\}}{\mathbb{P}(D=s,G=\mathrm{obs})}\right].$
Hence, $\theta_{o,s}$ can also be handled in the same manner as $\theta_{o}$.
In what follows, we examine the partial identification of $\theta_{o}$ based on experimental data alone and based on the combination of experimental and observational data.
2.3 Identified Sets for DTE Parameters under Experimental and Combined Data
We formally define the identified sets for the DTE parameter $\theta_{o}$ under two data scenarios: (i) experimental data alone and (ii) the combination of experimental and observational data. We begin with the case of experimental data only.
Let $\mathcal{F}^{\dagger}$ denote the class of CDFs for all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5.333Formally, $\mathcal{F}^{\dagger}$ is the set of all CDFs $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$. We begin by defining the identified set for the joint CDF $F^{\ast}$ under the experimental data as
$\mathcal{F}_{\mathrm{exp}}^{\ast}\equiv\left\{F\in\mathcal{F}^{\dagger}:F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp}),\ F_{X}(\cdot)=F_{X}^{\ast}(\cdot)\right\},$
where $F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ is the distribution of the observables $(Y,D,X)$ in the experimental data, and $F_{X}^{\ast}$ is the marginal distribution of $X$ in the entire population. Thus, $\mathcal{F}_{\mathrm{exp}}^{\ast}$ consists of all CDFs that satisfy the maintained assumptions and are consistent with the distribution of the observed experimental data and the marginal distribution of $X$ in the entire population.
We assume that $F_{X}^{\ast}$ is known, as this adjustment is necessary to account for potential imbalances in the covariate distributions between the experimental and observational data and to enable meaningful comparisons of the identified sets. When the two data sources are drawn from the same population (i.e., covariates are balanced), this assumption is unnecessary because $F_{X}^{\ast}$ can be identified from the experimental data as $F_{X}^{\ast}(\cdot)=F_{X|G}^{\ast}(\cdot\mid\mathrm{exp})$. In other cases, covariate information for the entire population is often available from external sources such as demographic datasets. Moreover, when the parameter of interest is the covariate-conditional effect $\theta_{o,x}$ or the effect $\theta_{o,\mathrm{exp}}$ defined for the experimental population, knowledge of $F_{X}^{\ast}$ is not required.
Given the identified set of CDFs $\mathcal{F}_{\mathrm{exp}}^{\ast}$, the identified set for $\theta_{o}$ under the experimental data is defined as
$\Theta_{I}\equiv\left\{\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]:F\in\mathcal{F}_{\mathrm{exp}}^{\ast}\right\}.$ (2)
This set consists of all parameter values that are attainable under some distribution in $\mathcal{F}_{\mathrm{exp}}^{\ast}$. Any parameter value outside $\Theta_{I}$ is incompatible with the experimental data or the maintained assumptions.
We next consider the case in which both experimental and observational data are available. We first define the identified set for the joint CDF $F^{\ast}$ under the combined data, analogous to $\mathcal{F}_{\mathrm{exp}}^{\ast}$, as follows:
$\mathcal{F}^{\ast}\equiv\left\{F\in\mathcal{F}^{\dagger}:F_{YDGX}=F_{YDGX}^{\ast}\right\},$
where $F_{YDGX}^{\ast}$ is the joint distribution of the observables $(Y,D,G,X)$ in the combined data. This set consists of all CDFs that satisfy the maintained assumptions and are consistent with the observed distribution $F_{YDGX}^{\ast}$. By construction, $\mathcal{F}^{\ast}\subseteq\mathcal{F}_{\mathrm{exp}}^{\ast}$, since the condition $F_{YDGX}=F_{YDGX}^{\ast}$ in $\mathcal{F}^{\ast}$ implies both $F_{YDX|G}(\cdot,\cdot,\cdot\mid\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot\mid\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$.
Given the identified set of CDFs $\mathcal{F}^{\ast}$, the identified set for $\theta_{o}$ under the combined data is defined as
$\Theta_{IC}\equiv\left\{\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]:F\in\mathcal{F}^{\ast}\right\}.$ (3)
This set consists of all parameter values that are attainable under some CDF in $\mathcal{F}^{\ast}$. Any value outside $\Theta_{IC}$ contradicts the maintained assumptions or the observed combined data.
Since $\mathcal{F}^{\ast}\subseteq\mathcal{F}_{\mathrm{exp}}^{\ast}$, it follows that $\Theta_{IC}\subseteq\Theta_{I}$, meaning that the identified set under the combined data is no larger than that under the experimental data alone. Our primary interest, however, lies in whether the strict inclusion $\Theta_{IC}\subset\Theta_{I}$ holds. This corresponds to asking whether combining experimental and observational data strictly tightens the identified set for $\theta_{o}$. This question is nontrivial because a strict inclusion $\mathcal{F}^{\ast}\subset\mathcal{F}_{\mathrm{exp}}^{\ast}$ at the level of CDFs does not necessarily imply a strict inclusion $\Theta_{IC}\subset\Theta_{I}$ at the level of the parameter. We examine this question in the following sections.
3 Characterization of the Identified Sets
To investigate whether the strict inclusion $\Theta_{IC}\subset\Theta_{I}$ holds, the definitions of $\Theta_{I}$ and $\Theta_{IC}$ in equations (2) and (3) are too abstract to provide direct insight. We therefore seek to characterize these identified sets in a more interpretable form. To this end, we employ the concept of bivariate copulas (Sklar, 1959), which serves as a central tool for the analysis in this and subsequent sections.
3.1 Characterization of $\Theta_{I}$ for Experimental Data
We begin with the case of using only experimental data. Under randomized treatment assignment (Assumption 2.3) and external validity (Assumption 2.4), the marginal distributions of the potential outcomes conditional on covariates, $(F_{Y_{1}|X}^{\ast},F_{Y_{0}|X}^{\ast})$, are identified as $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$ for d=0,1d=0,1.444This follows from the equalities $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y_{d}|GX}^{\ast}(\cdot|\mathrm{exp},x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$, where the first and second equalities follow from Assumptions 2.4 and 2.1, respectively. Since $F_{X}^{\ast}$ is assumed to be known, we can identify the joint distribution of each potential outcome and the covariates, $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$.
Let $\mathcal{C}$ denote the class of all bivariate copula functions. For each $x\in\mathcal{X}$, let $C^{\ast}(\cdot,\cdot|x)\in\mathcal{C}$ denote the true conditional copula given $x$, which reproduces the true conditional joint distribution $F_{Y_{1}Y_{0}|X}^{\ast}(y_{1},y_{0}|x)$ from the marginals $(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x))$ as follows:
$F_{Y_{1}Y_{0}|X}^{\ast}(y_{1},y_{0}|x)=C^{\ast}\left(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x)|x\right),$
where the existence of such a copula function is guaranteed by Sklar’s theorem (e.g., Nelsen, 2006, Theorem 2.3.3). [^5] Sklar’s theorem (Sklar, 1959) states that any joint distribution $F_{Y_{1}Y_{0}}$ can be expressed as a copula function of its marginals: $F_{Y_{1}Y_{0}}(y_{1},y_{0})=C(F_{Y_{1}}(y_{1}),F_{Y_{0}}(y_{0}))$ for some copula function $C$. Conversely, given any marginal distributions $F_{Y_{1}}$ and $F_{Y_{0}}$, the function $C(F_{Y_{1}}(y_{1}),F_{Y_{0}}(y_{0}))$, for any copula $C$, defines a valid bivariate distribution with those marginals. Using the true conditional copula $C^{\ast}(\cdot,\cdot|x)$, the parameter $\theta_{o}$ can be expressed as
$\theta_{o}$ $=\mathbb{E}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})\right]$
$=\mathbb{E}\left[\int\int\psi(y_{1},y_{0})dC^{\ast}\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\middle|X\right)\right].$
The true copula $C^{\ast}(\cdot,\cdot|x)$ is unknown. However, by replacing $C^{\ast}(\cdot,\cdot|x)$ with all possible copula functions $C(\cdot,\cdot|x)\in\mathcal{C}$, we obtain the identified set for $\theta_{o}$ based on the knowledge of $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$ as
$\widetilde{\Theta}_{I}\equiv$ $\left\{\theta:\theta=\mathbb{E}_{F_{X}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)|X)\right]\right$.
[…]
By Sklar’s theorem, the collection $\bigl\{C\bigl(F_{Y_{1}|X}^{\ast}(\cdot|x),F_{Y_{0}|X}^{\ast}(\cdot|x)\big|x\bigr):C(\cdot,\cdot|x)\in\mathcal{C}\bigr\}$ coincides with the set of all conditional joint CDFs of $(Y_{1},Y_{0})$ given $X=x$ that share the marginals $\bigl(F_{Y_{1}|X}^{\ast}(\cdot|x),F_{Y_{0}|X}^{\ast}(\cdot|x)\bigr)$.
The following proposition formally establishes that the identified set $\widetilde{\Theta}_{I}$, based on the distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$, coincides with the identified set $\Theta_{I}$ under the experimental data.
Proposition 3.1 (Characterization of $\Theta_{I}$).
The identified set $\Theta_{I}$ under the experimental data is equivalent to the identified set $\widetilde{\Theta}_{I}$ based on the distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$; that is, $\Theta_{I}=\widetilde{\Theta}_{I}$.
This result formally confirms that $\Theta_{I}$ is fully characterized by the marginal distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$. [^6] While many studies refer to $\widetilde{\Theta}_{I}$ as the identified set under experimental data, Proposition 3.1 provides the formal justification for this equivalence.
3.2 Characterization of $\Theta_{IC}$ for Combined Data
We next seek to characterize the identified set $\Theta_{IC}$ under the combined data. Given that the experimental data identify $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$, our starting point is to examine what additional information can be obtained by incorporating the observational data. The following lemma addresses this question.
Lemma 3.2 (Identification of $F_{Y_{1}SX}^{\ast}$ and $F_{Y_{0}SX}^{\ast}$).
Suppose that Assumptions 2.12.5 hold. Then, for any $(d,s)\in\{0,1\}^{2}$ and almost all $x\in\mathcal{X}$, both $\mathbb{P}(S=s|X=x)$ and $F_{Y_{d}|SX}^{\ast}(\cdot|s,x)$ are identified from the combined data $(Y,D,G,X)$ as follows:
$\mathbb{P}(S=s|X=x)=\mathbb{P}(D=s|G=\mathrm{obs},X=x)$ (4)
and
$F_{Y_{d}|SX}^{\ast}(\cdot|s,x)=\begin{cases}\begin{array}[]{c}F_{Y|DGX}^{\ast}(\cdot|s,\mathrm{obs},x)\\ \frac{F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)-\mathbb{P}(D=d|G=\mathrm{obs},X=x)\cdot F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)}{\mathbb{P}(D=s|G=\mathrm{obs},X=x)}\\ \end{array}&\begin{array}[]{c}\mbox{if }d=s\\ \mbox{otherwise}\\ \end{array}\end{cases}.$ (5)
Lemma 3.2 shows that combining experimental and observational data enables identification of the joint distributions $\left(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\right)$ of each potential outcome, self-selection variable, and covariates. Therefore, the combined data yield richer information than the experimental data alone, which identify only $\left(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast}\right)$.
In particular, $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)$ for $d\neq s$ is a counterfactual distribution and cannot be identified from either the experimental or observational data alone. Combining both data sources, however, makes it identifiable through the following decomposition: For $d\neq s$,
$F_{Y_{d}|X}^{\ast}(\cdot|x)=\mathbb{P}(S=d|X=x)\cdot F_{Y_{d}|SX}^{\ast}(\cdot|d,x)+\mathbb{P}(S=s|X=x)\cdot F_{Y_{d}|SX}^{\ast}(\cdot|s,x).$
Here, $F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)$ is identified from the experimental data, while $\mathbb{P}(S=\cdot\mid X=x)$ and $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)$ are identified from the observational data (see the proof for details). [^7] Long et al. (2008) also show a related result: the ATE $\mathbb{E}[Y_{1}-Y_{0}\mid S=s]$ for each self-selection group $s\in\{0,1\}$ can be identified by combining data from random assignment and self-selection sources.
Let $C^{\ast}(\cdot,\cdot|s,x)$ denote the true conditional copula given $S=s$ and $X=x$; that is,
$F_{Y_{1}Y_{0}|SX}^{\ast}(y_{1},y_{0}|s,x)=C^{\ast}\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x)\big|s,x\bigr).$
Then the true parameter value $\theta_{o}$ can be expressed as
$\theta_{o}$ $=\mathbb{E}_{F_{Y_{1}Y_{0}}^{\ast}}\left[\psi(Y_{1},Y_{0})\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|SX}^{\ast}(y_{1},y_{0}|S,X)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC^{\ast}\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right].$
Although the true conditional copula $C^{\ast}(\cdot,\cdot\mid s,x)$ is unknown, replacing it with all possible copulas $C(\cdot,\cdot|s,x)\in\mathcal{C}$ yields the identified set for $\theta_{o}$ based on $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$:
$\widetilde{\Theta}_{IC}\equiv$ $\left\{\theta:\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)|S,X)\right]\right$.
[…]
By Sklar’s theorem, for any $C\in\mathcal{C}$, the function $C\big(F_{Y_{1}|SX}^{\ast}(y_{1}\mid s,x),F_{Y_{0}|SX}^{\ast}(y_{0}\mid s,x)\mid s,x\big)$ defines a valid conditional joint CDF of $(Y_{1},Y_{0})$ given $S=s$ and $X=x$, with marginals $F_{Y_{1}|SX}^{\ast}(\cdot\mid s,x)$ and $F_{Y_{0}|SX}^{\ast}(\cdot\mid s,x)$.
We now ask whether $\widetilde{\Theta}_{IC}$ coincides with the identified set $\Theta_{IC}$ under the combined data. This question can be rephrased as whether $\Theta_{IC}$ is fully characterized by the self-selection joint distributions $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$. This question is nontrivial because combining experimental and observational data might yield additional identifying information beyond these joint distributions.
The following theorem establishes an affirmative answer to this question: the identified set $\Theta_{IC}$ is fully characterized by $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$.
Theorem 3.3 (Characterization of $\Theta_{IC}$).
The identified set $\Theta_{IC}$ under the combined data is equivalent to the identified set $\widetilde{\Theta}_{IC}$ based on the distributions $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$; that is, $\Theta_{IC}=\widetilde{\Theta}_{IC}$.
In summary, the identified set $\Theta_{I}$ under the experimental data is fully characterized by $(F_{Y_{0}X}^{\ast},F_{Y_{1}X}^{\ast})$ (Proposition 3.1), whereas the identified set $\Theta_{IC}$ under the combined data is fully characterized by the self-selection joint distributions $(F_{Y_{0}SX}^{\ast},F_{Y_{1}SX}^{\ast})$ (Theorem 3.3). This comparison suggests that combining experimental and observational data yields additional identifying power through the dependence between potential outcomes $(Y_{1},Y_{0})$ and self-selection $S$. The following section examines this point in detail.
Remark 3.1 (Selection-on-observables).
The comparison between $\widetilde{\Theta}_{I}$ and $\widetilde{\Theta}_{IC}$ reveals that if self-selection $S$ is independent of the potential outcomes $(Y_{1},Y_{0})$ conditional on $X$, then $\Theta_{I}$ and $\Theta_{IC}$ coincide. This implies that when selection-on-observables, $(Y_{1},Y_{0})\perp\!\!\!\perp D\mid X$, holds in the observational data, the combination of the two data sources provides no additional identifying power for $\theta_{o}$. In contrast to much of the identification literature, where selection-on-observables is desirable in observational data, identification benefits from selection-on-unobservables in the present context.
4 Bounds Analysis
In this section, we derive sharp bounds for the DTE parameter $\theta_{o}$ under both experimental and combined data. Our analysis builds on Fan et al. (2017), who study the identifying power of covariates for DTEs. Extending their framework, we examine the identifying power of the self-selection variable $S$, which, as shown in Theorem 3.3, captures the additional identifying power of the combined data.
Following Fan et al. (2017), we consider two classes of DTE parameters, where $\psi$ is specified as (i) a super-modular function and (ii) a $\varphi$-indicator function. For each case, we derive closed-form expressions for the sharp bounds on $\theta_{o}$ under both experimental and combined data. We further establish, for each case, necessary and sufficient conditions under which combining the two data sources strictly tightens the identified set (i.e., $\Theta_{IC}\subset\Theta_{I}$).
4.1 Identified Sets for Super-modular Functions and the Identification Power of Combined Data
We begin with DTR parameters represented by super-modular and sub-modular functions.
Definition 4.1 (Super/sub-modular).
(i) A function $\psi(\cdot,\cdot)$ is super-modular if, for all $y_{1}\leq y_{1}^{\prime}$ and $y_{0}\leq y_{0}^{\prime}$, $\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})\geq 0$. It is sub-modular if $\psi(\cdot,\cdot)$ is super-modular. (ii) A function $\psi(\cdot,\cdot)$ is strict super-modular if, for all $y_{1}<y_{1}^{\prime}$ and $y_{0}<y_{0}^{\prime}$, $\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})>0$. It is strict sub-modular if $\psi(\cdot,\cdot)$ is strict super-modular.
The functions $\psi(\cdot,\cdot)$ in Examples 2.3, 2.4, and 2.5 are either super-modular or sub-modular and thus fall within this framework. In particular, the function $\psi$ in Example 2.5 is strict super-modular. Cambanis et al. (1976) provide many other examples of super-modular and sub-modular functions; see also Fan et al. (2017) for further examples of strict super-modular and sub-modular functions.
We characterize $\Theta_{I}$ following the analysis of Fan et al. (2017). Define
$F^{\ast,(-)}(y_{1},y_{0})\equiv\mathbb{E}\left[M\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\right)\right]\mbox{\ and\ }$
$F^{\ast,(+)}(y_{1},y_{0})\equiv\mathbb{E}\left[W\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\right)\right],$
where $M(u,v)\equiv\max(u+v-1,0)$ and $W(u,v)\equiv\min(u,v)$ are the Fréchet–Hoeffding lower and upper bounds, respectively, for a bivariate distribution with marginals $(u,v)$.
When $\psi$ is super-modular, Theorem 3.2 of Fan et al. (2017) shows that under certain regularity conditions, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{1}X}^{\ast})$ is the interval $\left[\theta^{L},\theta^{U}\right]$, where
$\theta^{L}$ $\equiv\mathbb{E}_{F^{\ast,(-)}}[\psi(Y_{1},Y_{0})]=\mathbb{E}_{F_{X}^{\ast}}\left[\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|X),F_{Y_{0}|X}^{\ast,-1}(1-u|X))du\right]\mbox{\ and}$ (6)
$\theta^{U}$ $\equiv\mathbb{E}_{F^{\ast,(+)}}[\psi(Y_{1},Y_{0})]=\mathbb{E}_{F_{X}^{\ast}}\left[\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|X),F_{Y_{0}|X}^{\ast,-1}(u|X))du\right],$ (7)
with $F_{Y_{d}|X}^{\ast,-1}(u|x)\equiv\inf\bigl\{y:F_{Y_{d}|X}^{\ast,-1}(y|x)\geq u\bigr\}$ denoting the quintile function of $Y_{d}$ conditional on X=xX=x. [^8] These analytical expressions highlight the identification power of the covariates $X$, as shown by Fan et al. (2017).
Therefore, by Proposition 3.1, the identified set $\Theta_{I}$ under the experimental data is given by the interval $\left[\theta^{L},\theta^{U}\right]$. Fan et al. (2017) also propose inference methods based on the analytical expressions (6) and (7) for the sharp bounds.
We next characterize the identified set $\Theta_{IC}$ under the combined data. As in the previous case, we define
$F_{(-)}^{\ast}(y_{1},y_{0})\equiv\mathbb{E}\left[M\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\right)\right]\mbox{\ and\ }$
$F_{(+)}^{\ast}(y_{1},y_{0})\equiv\mathbb{E}\left[W\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\right)\right],$
where $F_{(\star)}^{\ast}(y_{1},y_{0})$, for $\star\in\{-,+\}$, differs from $F^{\ast,(\star)}(y_{1},y_{0})$ by the inclusion of the self-selection variable $S$.
Then, by applying Theorem 3.2 of Fan et al. (2017), we obtain that when $\psi$ is super-modular and certain regularity conditions hold, the identified set $\widetilde{\Theta}_{IC}$ based on $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$ is the interval $\left[\theta_{L},\theta_{U}\right]$, where
$\theta_{L}$ $\equiv\mathbb{E}_{F_{(-)}^{\ast}}[\psi(Y_{1},Y_{0})]=\mathbb{E}_{F_{SX}^{\ast}}\left[\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|S,X),F_{Y_{0}|SX}^{\ast,-1}(1-u|S,X))du\right]\mbox{\ and}$ (8)
$\theta_{U}$ $\equiv\mathbb{E}_{F_{(+)}^{\ast}}[\psi(Y_{1},Y_{0})]=\mathbb{E}_{F_{SX}^{\ast}}\left[\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|S,X),F_{Y_{0}|SX}^{\ast,-1}(u|S,X))du\right],$ (9)
with $F_{Y_{d}|SX}^{\ast,-1}(u|s,x)\equiv\inf\{y:F_{Y_{d}|SX}^{\ast,-1}(y|s,x)\geq u\}$ denoting the quintile function of $Y_{d}$ conditional on $S=s$ and $X=x$.
Therefore, by Theorem 3.3, the identified set $\Theta_{IC}$ under the combined data is given by the interval $\left[\theta_{L},\theta_{U}\right]$. We formalize these results in the following proposition.
Proposition 4.1 (Identified Sets for Super-modular Functions).
Let $\psi(y_{1},y_{0})$ be a super-modular and right-continuous function.
  • (i)
    Suppose that $\theta^{L}$ and $\theta^{U}$ exist (possibly infinite), and that either of the following conditions holds: (a) $\psi(y_{1},y_{0})$ is symmetric and both $\mathbb{E}[\psi(Y_{1},Y_{1})]$ and $\mathbb{E}[\psi(Y_{0},Y_{0})]$ are finite; (b) there exist some constants $\bar{y}_{0}$ and $\bar{y}_{1}$ such that $\mathbb{E}[\psi(Y_{1},\bar{y}_{0})]$ and $\mathbb{E}[\psi(\bar{y}_{1},Y_{0})]$ are finite, and at least one of $\theta^{L}$ and $\theta^{U}$ is finite. Then, $\Theta_{I}=\left[\theta^{L},\theta^{U}\right]$.
  • (ii)
    Suppose that $\theta_{L}$ and $\theta_{U}$ exist (possibly infinite), and that either conditions (a) or (b) holds, with $\theta^{L}$ and $\theta^{U}$ replaced by $\theta_{L}$ and $\theta_{U}$ in condition (b). Then, $\Theta_{IC}=\left[\theta_{L},\theta_{U}\right]$.
Proposition 4.1(ii) is our novel contribution, deriving the sharp bounds for $\theta_{o}$ under the combined experimental and observational data when $\psi$ is super-modular. Together with Lemma 3.2, this result enables computation of the sharp bounds via equations (8) and (9), in conjunction with (4) and (5). It also provides a basis for constructing inference procedures. [^9] Pointwise valid confidence sets for $\theta_{o}$ can be constructed by extending the inference procedure of Fan et al. (2017, Appendix B) to settings with nonparametric estimators of $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$, which can be obtained using Lemma 3.2.
The key difference between the sharp bounds $\theta^{L}$ ($\theta^{U}$) and $\theta_{L}$ ($\theta_{U}$) under the two data scenarios lies in the inclusion of the self-selection variable $S$ in the conditioning sets (see equations (8) and (9)). This inclusion implies that self-selection can offer additional identifying power for $\theta_{o}$ when experimental data are combined with observational data. In particular, $S$ may carry information about the dependence structure between $Y_{1}$ and $Y_{0}$ (for example, as in the Roy selection model), thereby tightening the identified set.
To see this formally, note that $\Theta_{I}$ is the set of values of $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ as $F$ ranges over all joint CDFs $F_{Y_{1}Y_{0}}$ with fixed marginals $F^{\ast,(-)}$ and $F^{\ast,(+)}$ satisfying $F^{\ast,(-)}(y_{1},y_{0})\leq F_{Y_{1}Y_{0}}(y_{1},y_{0})\leq F^{\ast,(+)}(y_{1},y_{0})$, whereas $\Theta_{IC}$ is the set of values of $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ as $F$ ranges over all joint CDFs $F_{Y_{1}Y_{0}}$ with fixed marginals $F_{(-)}^{\ast}$ and $F_{(+)}^{\ast}$ satisfying $F_{(-)}^{\ast}(y_{1},y_{0})\leq F_{Y_{1}Y_{0}}(y_{1},y_{0})\leq F_{(+)}^{\ast}(y_{1},y_{0})$. By Jensen’s inequality, it follows that
$F^{\ast,(-)}(y_{1},y_{0})$ $=\mathbb{E}\left[\max\left\{\mathbb{E}\left[F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1\middle|X\right],0\right\}\right]$
$\leq\mathbb{E}\left[\max\left\{F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1,0\right\}\right]$
$=F_{(-)}^{\ast}(y_{1},y_{0}),$
and similarly $F^{\ast,(+)}(y_{1},y_{0})\geq F_{(+)}^{\ast}(y_{1},y_{0})$. Since $F^{\ast,(-)}(\cdot,\cdot)\leq F_{(-)}^{\ast}(\cdot,\cdot)$ and $F_{(+)}^{\ast}(\cdot,\cdot)\leq F^{\ast,(+)}(\cdot,\cdot)$, we obtain $\theta^{L}\leq\theta_{L}$ and $\theta_{U}\leq\theta^{U}$, which implies that the data combination potentially shrinks the identified set.
For strict modular functions $\psi$, Theorem 4.2 below establishes the necessary and sufficient condition under which $\Theta_{IC}=\Theta_{I}$; otherwise, combining the two data sources yields a strictly smaller identified set (i.e., $\Theta_{IC}\subsetneq\Theta_{I}$).
Theorem 4.2 (Identification Power of Data Combination).
Let $\psi(y_{1},y_{0})$ be a strict super-modular and right-continuous function. Suppose that the four expectations in equations (6)–(9) exist (even if infinite valued), and that either conditions (a) and (b) in Proposition 4.1 (i) hold. Then $\Theta_{IC}=\Theta_{I}$ if and only if the following two conditions hold for $\psi_{c}$-almost all $(y_{1},y_{0})$: [^10] If $\psi(\cdot,\cdot)$ is super-modular and right continuous, it uniquely determines a non-negative measure $\psi_{c}$ on the Borel subsets of $\mathbb{R}^{2}$ such that for all $y_{1}\leq y_{1}^{\prime}$ and $y_{0}\leq y_{0}^{\prime}$, $\psi_{c}((y_{1},y_{1}^{\prime}]\times(y_{0},y_{0}^{\prime}])=\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})$. See Cambanis et al. (1976) and Rachev and Rüschendorf (2006).
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\middle|X\right)\in\{0,1\}\text{ a.s. }\mbox{\ and\ }$ (10)
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\middle|X\right)\in\{0,1\}\text{ a.s.}$ (11)
The conditions in equations (10) and (11) require that, conditional on $X$, the relative ordering of the conditional distributions $F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)$ and $F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)$ is degenerate in the sense that it does not vary with the self-selection variable $S$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Such conditions are highly restrictive and unlikely to hold when self-selection $S$ depends on the potential outcomes $Y_{1}$ and $Y_{0}$. In such cases, the conditional distributions $F_{Y_{1}|S,X}^{\ast}$ and $F_{Y_{0}|S,X}^{\ast}$ typically vary with $S$ in a way that violates these conditions. Therefore, when self-selection is outcome-dependent, combining experimental and observational data is likely to tighten the identified set for $\theta_{o}$.
A notable exception arises when self-selection is independent of the potential outcomes conditional on $X$ (i.e., when selection-on-observables holds). In this case, equations (10) and (11) are satisfied, and combining the two data sources provides no additional identifying power.
4.2 Identified Sets for $\varphi$-indicator Functions and the Identification Power of Data Combination
We now turn to the DTE parameters characterized by $\varphi$-indicator functions.
Definition 4.2 ($\varphi$-Indicator Functions).
Let $\varphi$ be a measurable function and define $\psi(y_{1},y_{0})\equiv\mathbf{1}\{\varphi(y_{1},y_{0})\leq\delta\}$ for a fixed $\delta$, where $\varphi(\cdot,\cdot)$ is monotone in each argument. We refer to this class of functions $\psi$ as the class of $\varphi$-indicator functions.
An important parameter in this class is the distribution function of treatment effect $F_{Y_{1}-Y_{0}}^{\ast}(\delta)$ (Example 2.2), which corresponds to the choice $\varphi(y_{1},y_{0})=y_{1}-y_{0}$. As a special case, the fraction of positive treatment effect (Example 2.1) is given by $\mathbb{P}(Y_{1}>Y_{0})=1-F_{Y_{1}-Y_{0}}^{\ast}(0)$.
We begin by considering the identified set $\Theta_{I}$ under the experimental data, or equivalently, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$. Let $\mathcal{Y}_{1}(x)$ and $\mathcal{Y}_{0}(x)$ denote the supports of $Y_{1}$ and $Y_{0}$ given $X=x$, respectively. Define
$F_{\min,\varphi}(\delta|x)$ $=\sup_{y\in\mathcal{Y}_{1}(x)}\max\left\{F_{Y_{1}|X}^{\ast}(y|x)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|x)-1,0\right\}\mbox{\ and}$
$F_{\max,\varphi}(\delta|x)$ $=1+\inf_{y\in\mathcal{Y}_{0}(x)}\min\left\{F_{Y_{1}|X}^{\ast}(y|x)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|x)-1,0\right\},$
where $\tilde{\varphi}_{y}(\delta|x)=\sup\left\{y_{0}\in\mathcal{Y}_{0}(x):\varphi(y,y_{0})<\delta\right\}$. For a fixed $\delta$, the set $\left\{y_{0}\in\mathcal{Y}_{0}(x):\varphi(y,y_{0})<\delta\right\}$ may be empty for some $y\in\mathcal{Y}_{1}(x)$ and $x\in\mathcal{X}$. In such cases, we define $\tilde{\varphi}_{y}(\delta|x)$ as minus infinity.
Fan et al. (2017) show that when $\psi$ is a $\varphi$-indicator function, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$ corresponds to the interval $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$, where $F_{\varphi}^{L}(\delta)=\mathbb{E}[F_{\min,\varphi}(\delta|X)]$ and $F_{\varphi}^{U}(\delta)=\mathbb{E}[F_{\max,\varphi}(\delta|X)]$. Proposition 3.1 therefore implies that the identified set $\Theta_{I}$ under the experimental data is also given by $\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$. Fan et al. (2017) also examine the identification power of the covariates $X$ through these analytical expressions for the sharp bounds.
We can similarly characterize the identified set $\widetilde{\Theta}_{IC}$ based on $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$. Let $\mathcal{Y}_{1}(s,x)$ and $\mathcal{Y}_{0}(s,x)$ denote the supports of $Y_{1}$ and $Y_{0}$ given $(S,X)=(s,x)$, respectively. Then $\widetilde{\Theta}_{IC}$ corresponds to the interval $\left[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)\right]$, where $F_{L,\varphi}(\delta)=\mathbb{E}[F_{\min,\varphi}(\delta|S,X)]$ and $F_{U,\varphi}(\delta)=\mathbb{E}[F_{\max,\varphi}(\delta|S,X)]$ with
$F_{\min,\varphi}(\delta|s,x)$ $=\sup_{y\in\mathcal{Y}_{1}(s,x)}\max\left\{F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1,0\right\},$ (12)
$F_{\max,\varphi}(\delta|s,x)$ $=1+\inf_{y\in\mathcal{Y}_{0}(s,x)}\min\left\{F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1,0\right\},$ (13)
and $\tilde{\varphi}_{y}(\delta|s,x)=\sup\left\{y_{0}\in\mathcal{Y}_{0}(s,x):\varphi(y,y_{0})<\delta\right\}$.
Therefore, by Theorem 3.3, the identified set $\Theta_{IC}$ under the combined data is given by the interval $\left[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)\right]$ when $\theta_{o}$ is represented by a $\varphi$-indicator function. The key difference from the experimental-data case lies in the inclusion of the self-selection variable $S$ in equations (12) and (13). Since $S$ may encode information about the dependence structure between $Y_{1}$ and $Y_{0}$, its inclusion can lead to a tighter identified set.
The following proposition summarizes these characterization results.
Proposition 4.3 (Identified Sets for $\varphi$-Indicator Functions).
Suppose that $\psi$ is a $\varphi$-indicator function and that $\varphi$ is continuous and non-decreasing in each argument.
  • (i)
    Suppose that $\mathcal{Y}_{1}(X)$ and $\mathcal{Y}_{0}(X)$ are Borel sets generated by intervals with measurable endpoints. Then $\Theta_{I}=\left[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)\right]$.
  • (ii)
    Suppose that $\mathcal{Y}_{1}(S,X)$ and $\mathcal{Y}_{0}(S,X)$ are Borel sets generated by intervals with measurable endpoints. Then $\Theta_{IC}=\left[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)\right]$.
Proposition 4.3(ii) is our novel contribution, deriving the sharp bounds for $\theta_{o}$—when represented by a $\varphi$-indicator function—under the combined experimental and observational data. Together with Lemma 3.2, this result enables direct computation of the identified set from the combined data.
Remark 4.1 (Numerical Example).
To illustrate the identifying power of the combined data, we present simple numerical examples in Appendix A. Focusing on $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, these examples demonstrate that the combined data can substantially narrow the identified set—and may even yield point identification—when self-selection depends on the potential outcomes.
Similar to Theorem 4.2, for a $\varphi$-indicator function we can establish sufficient and necessary conditions under which $\Theta_{IC}$ is a proper subset of $\Theta_{I}$. To simplify the technical argument, the following theorem presents this result for the case where $\mathcal{Y}_{d}(s,x)=\mathcal{Y}_{d}$ for $d=0,1$ and all $(s,x)\in\{0,1\}\times\mathcal{X}$.
Theorem 4.4 (Identification Power of Data Combination).
Suppose that $\psi$ is a $\varphi$-indicator function with $\varphi$ being continuous and non-decreasing in each argument, and that the condition in Proposition 4.3(ii) holds. Suppose further that $\mathcal{Y}_{d}(s,x)=\mathcal{Y}_{d}$ for $d=0,1$ and all $(s,x)\in\{0,1\}\times\mathcal{X}$. Then $\Theta_{IC}=\Theta_{I}$ if and only if, for almost all $x\in\mathcal{X}$, there exist $\bar{y}(x)$ and $\underline{y}(x)$ such that, for both $s\in\{0,1\}$, the function
$y\mapsto[F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1]$ (14)
attains its maximum at $\bar{y}(x)$ and its minimum at $\underline{y}(x)$.
The condition in Theorem 4.4 requires that the locations at which the function (14) attains its maximum and minimum be invariant to the self-selection variable $S$. This invariance condition is easily violated when self-selection $S$ depends on the potential outcomes $(Y_{1},Y_{0})$, as outcome-dependent selection typically alters the relative ordering of the conditional distributions across different values of $S$. In such cases, the theorem implies that combining experimental and observational data tightens the identified set.
A sufficient condition for the “if and only if” statement in Theorem 4.4 is that the self-selection variable $S$ is independent of the potential outcomes $(Y_{1},Y_{0})$ conditional on $X$. Under this selection-on-observables assumption, combining the two data sources provides no additional identifying power.
5 Additional Restrictions and Computational Approach
As in the philosophy of partial identification analysis (Manski, 2003), additional model restrictions can, when plausible, further narrow the identified set for $\theta_{o}$. Such restrictions, however, may complicate the analysis, particularly by making it difficult to derive sharp bounds. In this section, we introduce two restrictions that are plausible in many empirical contexts and present a computational approach to obtain the identified set under these restrictions with combined data.
5.1 Additional Restrictions
The first restriction is a form of positive dependence between the potential outcomes, specifically the mutually left-tail decreasing (LTD) condition (Joe, 2014). The potential outcomes $Y_{1}$ and $Y_{0}$ are said to be mutually LTD if they satisfy the following condition.
Assumption 5.1 (Positive Dependence).
The conditional distributions $\mathbb{P}(Y_{1}\leq t\mid Y_{0}\leq y,S=s,X=x)$ and $\mathbb{P}(Y_{0}\leq t\mid Y_{1}\leq y,S=s,X=x)$ are each non-increasing in $y$ almost everywhere, for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$. [^11] Conditioning on $S$ and $X$ may render this assumption restrictive. Alternatively, one may impose the same condition without conditioning on $S$ and $X$, resulting in a weaker assumption with less identifying power.
This assumption implies that individuals with higher potential outcomes under one treatment state are more likely to have higher potential outcomes under the other state. Such an assumption is plausible in many empirical contexts. For example, in a small-class-size program, students who perform well academically in either small- or regular-sized classes are also likely to perform well in the alternative class size. Frandsen and Lefgren (2021) consider a slightly stronger assumption and demonstrate that it can substantially tighten the identified set for the distribution of treatment effects. [^12] Frandsen and Lefgren (2021) consider a mutually stochastically increasing assumption, which requires the conditional distributions $\mathbb{P}(Y_{1}\leq t\mid Y_{0}=y,S=s,X=x)$ and $\mathbb{P}(Y_{0}\leq t\mid Y_{1}=y,S=s,X=x)$ to be non-increasing in $y$. This assumption is stronger than Assumption 5.1 (see Joe, 1997, Theorem 2.3). The same or related assumptions are also employed by Chetty et al. (2017) in their empirical study of income mobility and by Cui and Han (2025) in policy learning with distributional welfare.
The second restriction concerns the self-selection mechanism for treatment.
Assumption 5.2 (Generalized Roy Model Selection).
The inequality $\mathbb{P}(Y_{1}-Y_{0}>c\mid S=1,X=x)\geq\mathbb{P}(Y_{1}-Y_{0}>c\mid S=0,X=x)$ holds for all $c$ in the support of $Y_{1}-Y_{0}$ and for almost all $x\in\mathcal{X}$. [^13] One may also impose this assumption without conditioning on $X$, resulting in a weaker restriction with less identifying power.
This assumption implies that individuals who self-select into treatment are more likely to experience higher treatment effects than those who self-select into no-treatment. We refer to this as the generalized Roy model selection assumption, as it is implied by the selection behavior in the generalized Roy model. Because this restriction pertains to the self-selection mechanism, it cannot be leveraged using experimental data alone. Thus, an additional benefit of incorporating observational data is the ability to exploit such behavioral restrictions on self-selection.
5.2 Computational Approach
We propose a computational approach to obtain the sharp bounds for $\theta_{o}$ under the combined data, incorporating either or both Assumptions 5.1 and 5.2. When these assumptions are imposed in addition to Assumptions 2.12.5, the sharp lower and upper bounds for $\theta_{o}$ can be obtained by solving the following minimization and maximization problems:
$\underset{F_{Y_{1}Y_{0}SX}\in\mathcal{F}_{Y_{1}Y_{0}SX}}{\inf/\sup}\ \int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)$ (15)
s.t. $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}\quad\text{for }d=0,1;$ (16)
$F_{Y_{d}|Y_{d^{\prime}}\leq y,S=s,X=x}(t)\geq F_{Y_{d}|Y_{d^{\prime}}\leq y^{\prime},S=s,X=x}(t)$ (17)
[…]
$\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=1,X=x)\geq\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=0,X=x)$ (18)
$\mbox{for all }c\in\mathbb{R}\mbox{ and for almost all }x.$
The constraint in (16) follows from Theorem 3.3, which shows that the identified set for $\theta_{o}$ is fully characterized by $F_{Y_{1}SX}^{\ast}$ and $F_{Y_{0}SX}^{\ast}$. The constraints in (17) and (18) are directly derived from Assumptions 5.1 and 5.2, respectively. Note that the optimization problem (15)–(18) is formulated without using the treatment variable $D$ or the data source indicator $G$, since the constraint in (16) already incorporates all identifying information conveyed by these variables (Theorem 3.3). This formulation reduces the computational burden by minimizing the number of optimization variables.
When all random variables are discrete, the optimization problem (15)–(18) reduces to a finite-dimensional linear program, for which efficient algorithms and solvers are available. [^14] Many empirical applications, however, involve continuous outcomes and covariates. A common practical approach is to discretize these continuous variables, though this may come at the potential cost of losing sharpness in identification. The details of this linear programming formulation are provided in Appendix B. Inference methods for linear program solutions, including those proposed by Fang et al. (2023) and Cho and Russell (2024), are applicable in this setting.
Let $\widetilde{\mathcal{F}}^{\dagger}$ denote the class of CDFs $F$ for all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5 and 5.15.2 with $F^{\ast}$ replaced by $F$. The identified set for $\theta_{o}$ under the combined data and these assumptions is defined as
$\Theta_{IC}^{\dagger}\equiv\big\{\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]:F\in\widetilde{\mathcal{F}}^{\ast}\big\}$,
where $\widetilde{\mathcal{F}}^{\ast}\equiv\bigl\{F\in\widetilde{\mathcal{F}}^{\dagger}:F_{YDGX}=F_{YDGX}^{\ast}\bigr\}$.
The following proposition shows that $\Theta_{IC}^{\dagger}$ corresponds to an interval whose lower and upper bounds are given by the solutions to the minimization and maximization problems (15)–(18).
Proposition 5.1.
Let $\theta_{L}^{\ast}$ and $\theta_{U}^{\ast}$ be the solutions of the minimization and maximization problems (15)-(18), respectively. Then, as long as $\widetilde{\mathcal{F}}^{\ast}$ is nonempty, the identified set $\Theta_{IC}^{\dagger}$ is equal to the interval $\theta_{L}^{\ast},\theta_{U}^{\ast}]$.
This result enables us to compute the identified set for $\theta_{o}$ under the additional restrictions (Assumptions 5.1 and 5.2) by solving the linear program (15)–(18). Although our analysis focuses on the case in which both assumptions are imposed jointly, sharp bounds can also be obtained from optimization problems of the same form when each assumption is imposed individually.
6 Empirical Illustration
We illustrate the proposed approach using data from the DRPT study of Gaines and Kuklinski (2011), collected in Illinois as part of the 2008 Cooperative Campaign Analysis Project. The study examines the effect of negative campaign advertisements on candidate evaluations during the 2008 U.S. presidential election.
The sample consists of 483 adult Illinois residents who were randomly assigned to one of three groups: treatment ($n=118$), no-treatment ($n=129$), and self-selection ($n=236$). Individuals in the treatment group were exposed to negative campaign advertisements (e-flyers) about John McCain and Barack Obama, whereas those in the no-treatment group were not. Participants in the self-selection group were allowed to choose whether to view the advertisements, with 90 out of 236 participants opting in.
The outcome variable $Y$ is each respondent’s feeling thermometer rating toward each candidate, measured on a scale from 0 (very unfavorable) to 100 (very favorable), with 50 indicating neutrality. The dataset includes a single categorical covariate, $X$, indicating respondents’ partisanship: Republican ($n=207$), Democrat ($n=233$), and Independent ($n=43$).
Let $Y_{d,\text{McCain}}$ and $Y_{d,\text{Obama}}$ denote the potential feeling thermometer ratings toward John McCain and Barack Obama, respectively, under treatment status $d\in\{0,1\}$, where $d=1$ indicates exposure to the negative campaign advertisements. Our parameter of interest is $\mathbb{P}(Y_{1,j}<Y_{0,j})$, which represents the proportion of individuals whose evaluation of candidate $j\in\{\text{McCain},\text{Obama}\}$ is negatively affected by the campaign material. This parameter is particularly appealing because it remains well-defined and substantively meaningful even when the outcome $Y_{d,j}$ is ordinal rather than cardinal. This feature is important in our application, as the feeling thermometer rating reflects individuals’ subjective assessments that may lack a consistent cardinal interpretation (see, e.g., Wilcox et al. (1989) for discussion). In such a context, common causal parameters, such as the ATE $\mathbb{E}[Y_{1,j}-Y_{0,j}]$, may fail to provide a meaningful interpretation.
For each candidate $j\in\{\text{McCain},\text{Obama}\}$, we estimate the identified set of $\mathbb{P}(Y_{1,j}<Y_{0,j})$ both with and without inclusion of the self-selection sample ($G=\mathrm{obs}$) and with and without imposing Assumption 5.1 (Positive Dependence). Assumption 5.1 is motivated by the idea that individuals who hold a more favorable view of a candidate in the control state are likely to maintain relatively favorable views even when exposed to negative campaign advertisements, inducing positive dependence between the potential outcomes. We do not impose Assumption 5.2, as the self-selection of negative campaign advertisements is unlikely to follow the Generalized Roy selection model. Without Assumption 5.1, the identified sets are estimated using Proposition 4.3 (with $\varphi(y_{1},y_{0})=y_{1}-y_{0}$ and $\delta=0$), together with Lemma 3.2, replacing $F_{YDGX}^{\ast}$ with its empirical distribution. When Assumption 5.1 is imposed, the identified sets are estimated via linear programming using empirical distributions (see Appendix B for details).
[Table 1 about here]
Table 1 reports the estimated identified sets of $\mathbb{P}(Y_{1}<Y_{0})$ for each candidate, for the full population and specified subpopulations (Democrats, Republicans, and each self-selection type $s\in\{0,1\}$), across the four combinations of (i) including vs. excluding the self-selection sample ($G=\mathrm{obs}$) and (ii) imposing vs. not imposing Assumption 5.1. The estimated identified sets based solely on the random-assignment sample ($G=\mathrm{exp}$) are wide and therefore not particularly informative. By contrast, incorporating the self-selection sample markedly tightens the identified sets—by 22% and 79% for John McCain without and with Assumption 5.1, respectively, and by 24% and 67% for Barack Obama without and with Assumption 5.1.
In particular, the joint use of the self-selection sample and Assumption 5.1 (our baseline setup) yields sufficiently informative estimates. For John McCain, the baseline results suggest that at least 40% of individuals are negatively affected by the negative campaign advertisement, while at least 47% appear resistant. Qualitatively similar patterns are observed for Barack Obama (see Table 1). These findings point to a substantial yet highly heterogeneous impact of the negative advertisements.
Overall, these findings demonstrate the empirical value of combining random-assignment and self-selection data sources. They highlight a novel advantage of DRPT designs: improving the informativeness of DTE analyses through the incorporation of self-selection data.
7 Conclusion
This study investigates how combining experimental and observational data can improve the identification of DTE parameters. We show that the identified set under the combined data is fully characterized by the joint distribution of each potential outcome, latent self-selection variable, and covariates, with latent self-selection serving as the key source of additional identification power. For a broad class of DTE parameters represented by super/sub-modular functions and $\varphi$-indicator functions, we derive sharp bounds under the combined data. We further establish necessary and sufficient conditions under which the data combination strictly shrinks the identified set, suggesting that such shrinkage generally occurs unless selection-on-observables holds in the observational data. We also propose a linear programming approach for computing sharp bounds while incorporating additional structural assumptions, such as positive dependence of potential outcomes and generalized Roy model selection. An empirical application using DRPT data on negative campaign advertisements in the U.S. presidential election illustrates the practical value of combining random-assignment and self-selection data, and highlights a novel advantage of DRPT designs.
Table
Table 1: Estimated Identified Sets of $\mathbb{P}(Y_{1}<Y_{0})$ for John McCain and Barack Obama
Panel (a): John McCain
Inclusion of Population
Self-selection Asm.PD All Democrats Republicans Selected T Selected NT
No No $[0.13,0.90]$ $[0.12,0.88]$ $[0.11,0.92]$
Yes No $[0.21,0.81]$ $[0.20,0.81]$ $[0.20,0.84]$ $[0.36,0.88]$ $[0.11,0.76]$
No Yes $[0.17,0.79]$ $[0.16,0.79]$ $[0.15,0.80]$
Yes Yes $[0.40,0.53]$ $[0.33,0.56]$ $[0.45,0.60]$ $[0.46,0.56]$ $[0.36,0.52]$
Panel (b): Barack Obama
Inclusion of Population
Self-selection Asm.PD All Democrats Republicans Selected T Selected NT
No No $[0.17,0.89]$ $[0.12,0.88]$ $[0.17,0.87]$
Yes No $[0.22,0.77]$ $[0.13,0.73]$ $[0.27,0.81]$ $[0.30,0.65]$ $[0.18,0.84]$
No Yes $[0.18,0.72]$ $[0.12,0.74]$ $[0.17,0.70]$
Yes Yes $[0.38,0.56]$ $[0.43,0.57]$ $[0.26,0.56]$ $[0.35,0.43]$ $[0.41,0.63]$
  • Notes: Each panel reports the estimated identified sets of $\mathbb{P}(Y_{1}<Y_{0})$ for each candidate, covering the full population and specified subpopulations (Democrats, Republicans, and each self-selection type $s\in\{0,1\}$) across four combinations: (i) including vs. excluding the self-selection sample ($G=\mathrm{obs}$) and (ii) imposing vs. not imposing Assumption 5.1 (Positive Dependence). “Self-selection” indicates whether the self-selection sample is included, and “Asm. PD” indicates whether Assumption 5.1 is imposed. “Democrats” and “Republicans” refer to self-identified party affiliation. “Selected T” and “Selected NT” denote individuals who self-selected into viewing and not viewing the negative campaign advertisements, respectively.
Appendix
Appendix A Numerical Example
This appendix illustrates the identifying power gained by combining experimental and observational data through a simple numerical example. Suppose that the parameter of interest is $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$. Let $(Y_{1},Y_{0})$ be supported on $1,1]\times[-1,1]$ with the following piecewise uniform joint density:
$f_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})=\begin{cases}\begin{array}[]{c}0.0,\\ 0.5,\\ 0.5,\\ 0.0,\end{array}&\begin{array}[]{c}(y_{1},y_{0})\in[-1,0]\times[-1,0]\\ (y_{1},y_{0})\in(0,1]\times[-1,0]\\ (y_{1},y_{0})\in[-1,0]\times(0,1]\\ (y_{1},y_{0})\in(0,1]\times(0,1]\end{array}\end{cases},$ (19)
where $(Y_{1},Y_{0})$ is uniformly distributed within each subrectangle. Under this distribution, the true parameter value is $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})=0.5$. We do not introduce any covariate $X$.
Under the supposed distribution, the marginal distributions $F_{Y_{1}}^{\ast}$ and $F_{Y_{0}}^{\ast}$ are both uniform on [1,1][-1,1]. By Proposition 4.3, the sharp identified set for $\theta_{o}$ under the experimental data can be computed as $F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$ with $\varphi(y_{1},y_{0})=y_{0}-y_{1}$ and $\delta=0$, yielding $\Theta_{I}=[0,1]$. Hence, in this example, the experimental data alone provide no information about $\theta_{o}$.
We further suppose that $\mathbb{P}(S=1)=1/2$ and that the conditional joint densities of $(Y_{1},Y_{0})$ given $S=0,1$ are piecewise uniform as follows:
$f_{Y_{1}Y_{0}|S}^{\ast}(y_{1},y_{0}|1)$ $=a\cdot\mathbf{1}\{(y_{1},y_{0})\in(0,1]\times[-1,0]\}+b\cdot\mathbf{1}\{(y_{1},y_{0})\in[-1,0]\times(0,1]\};$
$f_{Y_{1}Y_{0}|S}^{\ast}(y_{1},y_{0}|0)$ $=b\cdot\mathbf{1}\{(y_{1},y_{0})\in(0,1]\times[-1,0]\}+a\cdot\mathbf{1}\{(y_{1},y_{0})\in[-1,0]\times(0,1]\},$
with $(a,b)=(0.8,0.2)$. These conditional densities imply that individuals who self-select into treatment are more likely to experience positive treatment effects, whereas those who self-select into no-treatment are more likely to experience negative treatment effects. Note that the assumed conditional densities $f_{Y_{1}Y_{0}|S}$ and the self-selection probability $\mathbb{P}(S=\cdot)$ are consistent with the unconditional density in (19) (i.e., $f_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})=\sum_{s=0,1}\mathbb{P}(S=s)\cdot f_{Y_{1}Y_{0}|S}^{\ast}(y_{1},y_{0}|s)$).
Under the assumed distribution, the conditional CDFs $F_{Y_{d}|S}^{\ast}(y_{1}|s)$ for $d=0,1$ and $s=0,1$ are given by
$F_{Y_{1}|S}^{\ast}(y_{1}|1)$ $=F_{Y_{0}|S}^{\ast}(y_{0}|0)=b\cdot(1+\min\{0,y\})+a\cdot\max\{y,0\};$
$F_{Y_{1}|S}^{\ast}(y_{1}|0)$ $=F_{Y_{0}|S}^{\ast}(y_{0}|1)=a\cdot(1+\min\{0,y_{1}\})+b\cdot\max\{y_{1},0\}.$
Together with $\mathbb{P}(S=1)=0.5$, the identified set $\Theta_{IC}$ under the combined data can be computed using Proposition 4.3(ii), yielding $\Theta_{IC}=[0.3,0.7]$. This interval is substantially narrower than $\Theta_{I}=[0,1]$, demonstrating that combining experimental and observational data can substantially improve the identification of $\theta_{o}$.
Now consider the case in which $a=1$ and $b=0$ in the conditional densities above. This corresponds to a Roy selection model, in which self-selection is determined by $S=\mathbf{1}\{Y_{1}-Y_{0}>0\}$. Under this specification, the conditional densities $f_{Y_{1}Y_{0}|S}^{\ast}$ and the self-selection probability $\mathbb{P}(S=\cdot)$ remain consistent with the unconditional density in (19). In this case, the identified set $\Theta_{IC}$, computed using Proposition 4.3(ii), collapses to the singleton $\Theta_{IC}=\{0.5\}$. Thus, the combined data achieve point identification, whereas the experimental data alone remain completely uninformative ($\Theta_{I}=[0,1]$). Notably, this point identification is attained without any prior knowledge of the Roy selection mechanism; rather, it arises solely from combining experimental and observational data.
Appendix B Linear Programming
This appendix shows that when all random variables are discrete, the optimization problems (15)–(18) can be formulated as a finite-dimensional linear program. For a generic distribution function $F_{Y_{1}Y_{0}SX}$, let $f_{Y_{1}Y_{0}SX}$ denote the density function. Let $f_{Y_{d}SX}^{\ast}$ denote the true density of $(Y_{d},S,X)$ induced by $F_{Y_{d}SX}^{\ast}$, which can be identified as in Lemma 3.2.
The optimization problems (15)–(18) can then be reformulated in terms of the density functions as follows:
$\underset{f_{Y_{1}Y_{0}SX}}{\min/\max}\ \sum_{(y_{1},y_{0},s,x)}\psi(y_{1},y_{0})f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)$ (20)
s.t. $\sum_{(y_{1},y_{0},s,x)}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)=1\mbox{ and }0\leq f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)\leq 1\ \forall y_{1},y_{0},s,x;$ (21)
$\sum_{y_{d}\leq y,y_{d^{\prime}}\leq+\infty,s^{\prime}\leq s,x^{\prime}\leq x}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s^{\prime},x^{\prime})=F_{Y_{d}SX}^{\ast}(y,s,x)\ \ \forall y,s,x,d,d^{\prime}\mbox{ with }d\neq d^{\prime};$ (22)
$\frac{\sum_{y_{d}\leq t,Y_{d^{\prime}}\leq y}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)}{\sum_{y_{d^{\prime}}\leq y}f_{Y_{d^{\prime}}SX}^{\ast}(y_{d^{\prime}},s,x)}-\frac{\sum_{y_{d}\leq t,y_{d^{\prime}}\leq y^{\prime}}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)}{\sum_{y_{d^{\prime}}\leq y^{\prime}}f_{Y_{d^{\prime}}SX}^{\ast}(y_{d^{\prime}},s,x)}\geq 0$ (23)
$\ \forall t,y,y^{\prime},d,d^{\prime},s,x\mbox{ with }y^{\prime}\geq y\mbox{ and }d\neq d^{\prime};$
$\frac{\sum_{y_{1}-y_{0}>c}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},1,x)}{f_{SX}^{\ast}(1,x)}-\frac{\sum_{y_{1}-y_{0}>c}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},0,x)}{f_{SX}^{\ast}(0,x)}\geq 0\ \ \forall c,x.$ (24)
The constraint in equation (21) ensures that $f_{Y_{1}Y_{0}SX}$ is a valid probability density function, while equations (22)–(24) correspond to the original constraints in (16)–(18), respectively. Note that the denominators in (23) and (24) involve the true densities $f_{Y_{d^{\prime}}SX}^{\ast}$ and $f_{SX}^{\ast}$, rather than $f_{Y_{d^{\prime}}SX}$ and $f_{SX}$. This substitution is valid because (22) implies $f_{Y_{d}SX}=f_{Y_{d}SX}^{\ast}$ for $d\in\{0,1\}$.
Since $f_{Y_{1}Y_{0}SX}$ takes finitely many values and all constraints in (21)–(24) are linear in $f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)$, the optimization problem (20)–(24) is a finite-dimensional linear program.
The following remark explains how we estimate the identified sets under Assumption 5.1 in the empirical application presented in Section 6.
Remark B.1 (Empirical Application in Section 6).
In our empirical application in Section 6, when Assumption 5.1 is imposed together with the self-selection sample, we estimate the sharp bounds by solving the linear program (20)–(23), setting $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}<y_{0}\}$ and replacing $F_{Y_{d}SX}^{\ast}$ and $f_{SX}^{\ast}$ with their empirical counterparts constructed according to Lemma 3.2.151515Specifically, each component in equations (4) and (5) is estimated using the corresponding empirical distribution or density. [^16] For equation (23), which is derived from Assumption 5.1, we exclude $t$- and $y$-values below the 2.5th percentile and above the 97.5th percentile of the estimated outcome distribution to mitigate boundary bias. To reduce memory usage and computation time, we additionally thin the grid by retaining every third $t$- and $y$-value. We do not impose the constraint in equation (24) because Assumption 5.2 is not maintained.
When Assumption 5.1 is imposed without the self-selection sample, we estimate the sharp bounds analogously using the corresponding linear program, but excluding the self-selection variable. Specifically, we drop the self-selection variable $s$ from the linear program (20)–(23) and replace $F_{Y_{d}X}^{\ast}$ and $f_{X}^{\ast}$ with their empirical counterparts based on the sample with $G=\mathrm{exp}$. [^17] That is, estimation relies solely on the experimental subsample.
We solve the linear programs using Gurobi. In finite samples, the optimization problem may become infeasible due to conflicting constraints. In such cases, we apply Gurobi’s feasrelax procedure to constraint (23), which introduces slack variables and solves an auxiliary optimization problem that minimizes the total relaxation required. This approach restores feasibility while keeping deviations from the original constraints as small as possible.
When estimating $\mathbb{P}(Y_{1}<Y_{0}\mid X=x)$, we restrict the sample to units with $X=x$. To estimate $\mathbb{P}(Y_{1}<Y_{0}\mid S=s)$, following the discussion in Section 2.2, we define $\psi(y_{1},y_{0})=\mathbf{1}\{y_{0}>y_{1}\}\cdot\mathbf{1}\{S=s\}/\mathbb{P}(S=s)$, where $\mathbb{P}(S=s)$ is estimated by the sample fraction of $D=s$ among units with $G=\mathrm{obs}$.
Appendix C Proofs and a Preliminary Lemma
Proof of Lemma 3.2.
We begin by showing equation (4). By Assumption 2.2,
$\mathbb{P}(D=s\mid G=\mathrm{obs},X=x)=\mathbb{P}(S=s\mid G=\mathrm{obs},X=x).$
Then, by Assumption 2.4, it follows that
$\mathbb{P}(S=s\mid G=\mathrm{obs},X=x)=\mathbb{P}(S=s\mid X=x).$
Combining these results yields equation (4).
We next consider the identification of $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)$, thereby showing equation (5). When $d=s$, we have
$F_{Y\mid DGX}^{\ast}(\cdot\mid d,\mathrm{obs},x)=F_{Y_{d}\mid DGX}^{\ast}(\cdot\mid d,\mathrm{obs},x)=F_{Y_{d}\mid SGX}^{\ast}(\cdot\mid s,\mathrm{obs},x)=F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x),$
where the second equality follows from Assumption 2.2, and the third from Assumption 2.4. Hence, equation (5) hols for $d=s$.
For the case $d\neq s$, we begin with the decomposition:
$F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)$ $=\mathbb{P}(S=d\mid X=x)\cdot F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)+\mathbb{P}(S=s\mid X=x)\cdot F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x).$
Rearranging yields
$F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)=\frac{F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)-\mathbb{P}(S=d\mid X=x)\cdot F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)}{\mathbb{P}(S=s\mid X=x)}.$
In this expression, $F_{Y_{d}|X}^{\ast}(\cdot|x)$ is identified as $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$ under Assumptions 2.3 and 2.4. Moreover, $\mathbb{P}(S=\cdot|X=x)$ and $F_{Y_{d}|SX}(\cdot|d,x)$ are identified as $\mathbb{P}(S=\cdot|X=x)=\mathbb{P}(D=\cdot|G=\mathrm{obs},X=x)$ and $F_{Y_{d}|SX}^{\ast}(\cdot|d,x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$, as shown above. Therefore, equation (5) holds for $d\neq s$. ∎
Lemma C.1.
For any distribution function $F\in\mathcal{F}^{\ast}$, the following hold:
  • (i)
    $F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|s,x)=F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|s,\mathrm{obs},x)$ for almost all $x$ and any $s=0,1$;
  • (ii)
    $F_{S|X}(\cdot|x)=F_{D|GX}(\cdot|G=\mathrm{obs},x)$ for almost all $x$.
Proof.
Fix any $F\in\mathcal{F}^{\ast}$. The result (i) follows from
$F_{Y_{1}Y_{0}\mid DGX}(\cdot,\cdot\mid d,\mathrm{obs},x)=F_{Y_{1}Y_{0}\mid SGX}(\cdot\mid d,\mathrm{obs},x)\ =F_{Y_{1}Y_{0}\mid SX}(\cdot\mid d,x),$
where the first equality follows from Assumption 2.2 and the second from Assumption 2.4. The result (ii) follows by the same argument as in the first part of the proof of Lemma 3.2. ∎
Proof of Proposition 3.1.
We begin by showing that $\Theta_{I}\subseteq\widetilde{\Theta}_{I}$. Fix any $\theta\in\Theta_{I}$. By definition, there exists a distribution function $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Since $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$) and $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$, it follows that for each $d\in\{0,1\}$ and all $x$,
$F_{Y_{d}|X}(\cdot|x)=F_{Y|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|X}^{\ast}(\cdot|x),$ (25)
where the first and third equalities follow from Assumptions 2.1, 2.3, and 2.4, and the second from $F_{YDX\mid G}(\cdot,\cdot,\cdot\mid\mathrm{exp})=F_{YDX\mid G}^{\ast}(\cdot,\cdot,\cdot\mid\mathrm{exp})$.
By Sklar’s Theorem (e.g., Nelsen, 2006, Theorem 2.3.3), there exists a conditional copula function $C(\cdot,\cdot\mid x)\in\mathcal{C}$ such that, for almost all $x$, the conditional joint distribution $F_{Y_{1}Y_{0}\mid X}$ can be written as
$F_{Y_{1}Y_{0}|X}(\cdot,\cdot|x)=C(F_{Y_{1}|X}(\cdot|x),F_{Y_{0}|X}(\cdot|x)|x).$
It then follows that
$\theta$ $=\mathbb{E}_{F_{Y_{1}Y_{0}}}\left[\psi(Y_{1},Y_{0})\right]$
$=\mathbb{E}_{F_{X}}\left[\mathbb{E}_{F_{Y_{1}Y_{0}|X}}\left[\psi(Y_{1},Y_{0})\middle|X\right]\right]$
$=\mathbb{E}_{F_{X}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|X}(y_{1},y_{0}|X)\right]$
$=\mathbb{E}_{F_{X}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|X}(y_{1}|X),F_{Y_{0}|X}(y_{0}|X)\big|X\bigr)\right]$
$=\mathbb{E}_{F_{X}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\big|X\bigr)\right]$
$\in\widetilde{\Theta}_{IC},$
where the fifth line follows from (25) and the condition $F_{X}=F_{X}^{\ast}$ in $\mathcal{F}_{\mathrm{exp}}^{\ast}$, and the last line follows from the definition of $\widetilde{\Theta}_{I}$. Hence, $\theta\in\widetilde{\Theta}_{I}$. Since this argument holds for any $\theta\in\Theta_{I}$, we conclude that $\Theta_{I}\subseteq\widetilde{\Theta}_{I}$.
We next show that $\widetilde{\Theta}_{I}\subseteq\Theta_{I}$. Fix any $\theta\in\widetilde{\Theta}_{I}$. By the definition of $\widetilde{\Theta}_{I}$, there exists a conditional copula function $C(\cdot,\cdot|x)\in\mathcal{C}$ such that
$\theta=\mathbb{E}_{F_{X}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\big|X\bigr)\right].$ (26)
We now show that there exists a distribution function $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ hierarchically, defined by
$F_{DGX}=F_{DGX}^{\ast};$ (27)
$F_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|D,G,X)=C\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\middle|X\right);$ (28)
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{obs},X)=\mathbf{1}\left\{D\leq s\right\};$ (29)
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{exp},X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X);$ (30)
$F_{Y|Y_{1}Y_{0}DSGX}(y)=\mathbf{1}\{DY_{1}+(1-D)Y_{0}\leq y\}.$ (31)
We will show that $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$) and (ii) $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$.
We begin with condition (i). From equation (31), we have $Y=DY_{1}+(1-D)Y_{0}$ a.s., and hence $F$ satisfies Assumption 2.1. Moreover, equation (29) implies $\mathbb{P}_{F}(S=D|G=\mathrm{obs},X)=1$ a.s., and hence $F$ satisfies Assumption 2.2.
Regarding Assumption 2.3, equation (28) implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ does not depend on $d$ (i.e., $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|1,\mathrm{exp},x)=F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|0,\mathrm{exp},x)$). Therefore, $F$ satisfies Assumption 2.3.
As for Assumption 2.4, by construction in equation (28), $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,g,x)$ does not depend on $d$ or $g$. Hence, $F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{exp},x)=F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x)$; that is, $F_{Y_{1}Y_{0}|GX}$ is invariant to $G$. Moreover, equations (29) and (30) imply that
$F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{exp},X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X),$
so $F_{S|Y_{1}Y_{0}GX}$ is also invariant to $G$. It then follows that, for almost all $(y_{1},y_{0},s,g,x)$,
$F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ $=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}GX}(s|u,v,g,x)dF_{Y_{1}Y_{0}|GX}(u,v|g,x)$
$=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}X}(s|u,v,x)dF_{Y_{1}Y_{0}|X}(u,v|x).$
Thus, $F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ does not depend on $g$, showing that $F$ satisfies Assumption 2.4.
The distribution $F$ also satisfies Assumption 2.5, since $F_{DGX}=F_{DGX}^{\ast}$ (equation (27)) and $F^{\ast}$ satisfies Assumption 2.5. Overall, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$).
We next consider condition (ii), namely that $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$. From equation (28), Sklar’s theorem implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ has conditional marginal distributions corresponding to $F_{Y_{1}|X}^{\ast}(\cdot|x)$ and $F_{Y_{0}|X}^{\ast}(\cdot|x)$. Hence, for almost all $d$ and $x$,
$F_{Y|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x),$
where the last equality follows from Assumptions 2.1, 2.3, and 2.4. Combining this with $F_{DGX}=F^{\ast}_{DGX}$ (equation (27)) yields $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$. Moreover, equation (27) leads to $F_{X}=F_{X}^{\ast}$. Hence, $F$ satisfies condition (ii).
We subsequently show that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. Note first that, by Assumptions 2.3 and 2.4, $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|X}(\cdot,\cdot|X)$. It then follows that
$\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ $=\mathbb{E}_{F_{X}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|X}(y_{1},y_{0}|X)\right]$
$=\mathbb{E}_{F_{X}}\left[\mathbb{E}_{F_{DG|X}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|D,G,X)\middle|X\right]\right]$
$=\mathbb{E}_{F_{X}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\middle|X\right)\right]$
$=\mathbb{E}_{F_{X}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\middle|X\right)\right]$
$=\theta,$
where the second line uses $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|X}(\cdot,\cdot|X)$, the third follows from equation (28), the fourth from equation (27), and the last from equation (26).
Consequently, we have shown that $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$, implying $\theta\in\Theta_{I}$. Since this argument holds for any $\theta\in\widetilde{\Theta}_{I}$, it follows that $\widetilde{\Theta}_{I}\subseteq\Theta_{I}$. ∎
Proof of Theorem 3.3.
We begin by showing that $\Theta_{IC}\subseteq\widetilde{\Theta}_{IC}$. Fix any $\theta\in\Theta_{IC}$. By definition, there exists a distribution function $F\in\mathcal{F}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Since $F_{YDGX}=F_{YDGX}^{\ast}$ holds and $F$ satisfies Assumptions 2.12.5, Lemma 3.2 implies that $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$ for $d=0,1$.
By Sklar’s theorem (e.g., Nelsen, 2006, Theorem 2.3.3), there exists a conditional copula function $C(\cdot,\cdot|s,x)\in\mathcal{C}$ such that, for almost all $(s,x)$,
$F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|s,x)=C(F_{Y_{1}|SX}(\cdot|s,x),F_{Y_{0}|SX}(\cdot|s,x)|s,x).$
Then it follows that
$\theta$ $=\mathbb{E}_{F_{Y_{1}Y_{0}}}\left[\psi(Y_{1},Y_{0})\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|SX}(y_{1},y_{0}|S,X)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}(y_{1}|S,X),F_{Y_{0}|SX}(y_{0}|S,X)\big|S,X\bigr)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right]$
$\in\widetilde{\Theta}_{IC},$
where the fourth line follows from $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, and the final inclusion follows from the definition of $\widetilde{\Theta}_{IC}$. Thus, $\theta\in\widetilde{\Theta}_{IC}$. Since this holds for any $\theta\in\Theta_{IC}$, we conclude that $\Theta_{IC}\subseteq\widetilde{\Theta}_{IC}$.
We next show that $\widetilde{\Theta}_{IC}\subseteq\Theta_{IC}$. Fix any $\theta\in\widetilde{\Theta}_{IC}$. By the definition of $\widetilde{\Theta}_{IC}$, there exists a conditional copula function $C(\cdot,\cdot|s,x)\in\mathcal{C}$ such that
$\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right].$ (32)
We will show that there exists a distribution function $F\in\mathcal{F}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ hierarchically, defined by
$F_{DGX}=F_{DGX}^{\ast};$ (33)
$F_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|D,\mathrm{obs},X)=C\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|D,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|D,\mathrm{obs},X)\middle|D,X\right);$ (34)
$F_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|GX}(y_{1},y_{0}|\mathrm{obs},X);$ (35)
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{obs},X)=\mathbf{1}\{D\leq s\};$ (36)
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{exp},X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X);$ (37)
$F_{Y|Y_{1}Y_{0}DSGX}(y|Y_{1},Y_{0},D,S,G,X)=\mathbf{1}\{DY_{1}+(1-D)Y_{0}\leq y\}.$ (38)
We will show that $F\in\mathcal{F}^{\ast}$ and that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$) and (ii) $F_{YDGX}=F_{YDGX}^{\ast}$.
We begin with condition (i). From equation (38), we have $Y=DY_{1}+(1-D)Y_{0}$ a.s. Hence, $F$ satisfies Assumption 2.1. Moreover, equation (36) implies $\mathbb{P}_{F}(S=D|G=\mathrm{obs},X)=1$ a.s. Hence, $F$ satisfies Assumption 2.2.
Regarding Assumption 2.3, equation (35) implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot\mid d,\mathrm{exp},x)$ does not depend on $d$. Therefore, $F$ satisfies Assumption 2.3.
As for Assumption 2.4, equation (35) implies that, for almost all $x$,
$F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{exp},x)$ $=\mathbb{E}_{F_{D}}[F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},x)]$
$=\mathbb{E}_{F_{D}}[F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x)]=F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x).$
Thus, $F_{Y_{1}Y_{0}|GX}$ does not depend on $G$. Moreover, equation (37) leads to
$F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\exp,X)$ $=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X),$
Hence, for almost all $(y_{1},y_{0},s,g,x)$,
$F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ $=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}GX}(s|u,v,g,x)dF_{Y_{1}Y_{0}|GX}(u,v|g,x)$
$=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}X}(s|u,v,x)dF_{Y_{1}Y_{0}|X}(u,v|x).$
Therefore, $F_{Y_{1}Y_{0}S\mid GX}(y_{1},y_{0},s\mid g,x)$ does not depend on $g$, and $F$ satisfies Assumption 2.4.
The distribution $F$ also satisfies Assumption 2.5, because $F_{DG\mid X}=F_{DG\mid X}^{\ast}$ and $F^{\ast}$ satisfies Assumption 2.5. To sum up, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$).
We now turn to condition (ii), namely that $F_{YDGX}=F_{YDGX}^{\ast}$. From equation (34), Sklar’s theorem implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{obs},x)$ has marginals equal to $F_{Y_{1}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$ and $F_{Y_{0}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$; that is, for almost all $d$ and $x$,
$F_{Y_{d}|DGX}(\cdot|d,\mathrm{obs},x)=F_{Y_{d}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x).$ (39)
Regarding $F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)$, we have for almost all $(d,x)$,
$F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)$ $=F_{Y_{d}|GX}(\cdot|\mathrm{obs},x)=F_{Y_{d}|GX}^{\ast}(\cdot|\mathrm{obs},x)$
$=F_{Y_{d}|GX}^{\ast}(\cdot|\mathrm{exp},x)=F_{Y_{d}|DGX}^{\ast}(\cdot|d,\mathrm{exp},x),$ (40)
where the first equality follows from equation (35), the second from equation (39), the third from Assumption 2.4, and the last from Assumption 2.3.
From equation (38), we have $F_{Y|DGX}(\cdot|d,g,x)=F_{Y_{d}|DGX}(\cdot|d,g,x)$ for almost all $(d,g,x)$. Together with results (39) and (40), this implies $F_{Y|DGX}(\cdot|d,g,x)=F_{Y|DGX}^{\ast}(\cdot|d,g,x)$ for almost all $(d,g,x)$. Combining this with $F_{DGX}=F_{DGX}^{\ast}$ (equation (33)) yields $F_{YDGX}=F_{YDGX}^{\ast}$. Hence, $F$ satisfies condition (ii).
We subsequently show that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. It follows that
$\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ $=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|SX}(y_{1},y_{0}|S,X)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|S,\mathrm{obs},X)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|S,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|S,\mathrm{obs},X)\middle|S,X\right)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|S,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|S,\mathrm{obs},X)\middle|S,X\right)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right]$
$=\theta,$
where the second and fifth equalities follow from Lemma C.1(i), the third from equation (34), and the last from equation (32). The fourth equality follows since
$F_{SX}(s,x)$ $=\int_{(-\infty,x]}F_{S|X}(s|u)dF_{X}(u)=\int_{(-\infty,x]}F_{D|GX}(s|\mathrm{obs},u)dF_{X}(u)$
$=\int_{(-\infty,x]}F_{D|GX}^{\ast}(s|\mathrm{obs},u)dF_{X}^{\ast}(u)=\int_{(-\infty,x]}F_{S|X}^{\ast}(s|u)dF_{X}^{\ast}(u)$
$=F_{SX}^{\ast}(s,x),$
where the second equality uses Lemma C.1(ii), the third equation uses (33), and the fourth again uses Lemma C.1(ii). Thus, we have established that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$.
Consequently, we have shown that $F\in\mathcal{F}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$, implying $\theta\in\Theta_{IC}$. Since this holds for any $\theta\in\widetilde{\Theta}_{IC}$, we conclude that $\widetilde{\Theta}_{IC}\subseteq\Theta_{IC}$. ∎
Proof of Proposition 4.1.
We first prove part (i). Under the stated conditions and by the definition of $\widetilde{\Theta}_{I}$, Theorem 3.2(i) in Fan et al. (2017) implies that $\widetilde{\Theta}_{I}=[\theta^{L},\theta^{U}]$. Proposition 3.1 then yields $\Theta_{I}=\widetilde{\Theta}_{I}=[\theta^{L},\theta^{U}]$.
We now turn to part (ii). Under the stated conditions and by the definition of $\widetilde{\Theta}_{IC}$, Theorem 3.2(i) in Fan et al. (2017), applied with $X$ replaced by $(S,X)$, implies that $\widetilde{\Theta}_{IC}=[\theta_{L},\theta_{U}]$. Theorem 3.3 then yields $\Theta_{IC}=\widetilde{\Theta}_{IC}=[\theta_{L},\theta_{U}]$. ∎
Proof of Theorem 4.2.
Recall the definitions of $F^{\ast,(-)}(y_{1},y_{0})$, $F^{\ast,(+)}(y_{1},y_{0})$, $F_{(-)}^{\ast}(y_{1},y_{0})$, and $F_{(+)}^{\ast}(y_{1},y_{0})$ in Section 4.1. For any $(y_{1},y_{0})$, by Jensen’s inequality, we obtain
$F^{\ast,(-)}(y_{1},y_{0})$ $=\mathbb{E}\left[\max\left\{F_{Y_{1}|X}^{\ast}(y_{1}|X)+F_{Y_{0}|X}^{\ast}(y_{0}|X)-1,0\right\}\right]$
$=\mathbb{E}\left[\max\left\{\mathbb{E}\left[F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1\middle|X\right],0\right\}\right]$
$\leq\mathbb{E}\left[\max\left\{F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1,0\right\}\right]$
$=F_{(-)}^{\ast}(y_{1},y_{0}),$ (41)
and
$F^{\ast,(+)}(y_{1},y_{0})$ $=\mathbb{E}\left[F_{Y_{0}|X}^{\ast}(y_{0}|X)+\min\left\{F_{Y_{1}|X}^{\ast}(y_{1}|X)-F_{Y_{0}|X}^{\ast}(y_{0}|X),0\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\middle|X\right]+\min\left\{\mathbb{E}\left[F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\middle|X\right],0\right\}\right]$
$\geq\mathbb{E}\left[F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)+\min\left\{F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X),0\right\}\right]$
$=F_{(+)}^{\ast}(y_{1},y_{0}).$ (42)
For any $s\in\{0,1\}$ and $x\in\mathcal{X}$, let
$\theta^{L}(x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|x),F_{Y_{0}|X}^{\ast,-1}(1-u|x))du$
$=\int\int\psi(y_{1},y_{0})dM(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x)),$
$\theta^{U}(x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|x),F_{Y_{0}|X}^{\ast,-1}(u|x))du$
$=\int\int\psi(y_{1},y_{0})dW(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x)),$
$\theta_{L}(s,x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|s,x),F_{Y_{0}|SX}^{\ast,-1}(1-u|s,x))du$
$=\int\int\psi(y_{1},y_{0})dM(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x)),$
and
$\theta_{U}(s,x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|s,x),F_{Y_{0}|SX}^{\ast,-1}(u|s,x))du$
$=\int\int\psi(y_{1},y_{0})dW(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x)).$
Then $\theta^{L}=\mathbb{E}[\theta^{L}(X)]$, $\theta^{U}=\mathbb{E}[\theta^{U}(X)]$, $\theta_{L}=\mathbb{E}[\theta_{L}(S,X)]$, and $\theta_{U}=\mathbb{E}[\theta_{U}(S,X)]$.
Under condition (a) of Proposition 4.1(i), it follows from equation (5) in Cambanis et al. (1976) that
$2\theta^{U}(X)$ $=\mathbb{E}[\psi(Y_{1},Y_{1})|X]+\mathbb{E}[\psi(Y_{0},Y_{0})|X]-\int\int A_{W}^{\ast}(X)d\psi_{c}(y_{1},y_{0}),$ (43)
$2\theta_{U}(S,X)$ $=\mathbb{E}[\psi(Y_{1},Y_{1})|S,X]+\mathbb{E}[\psi(Y_{0},Y_{0})|S,X]-\int\int A_{W}^{\ast}(S,X)d\psi_{c}(y_{1},y_{0}),$ (44)
where
$A_{W}^{\ast}(X)$ $\equiv F_{Y_{1}|X}^{\ast}(y_{1}\vee y_{0}|X)+F_{Y_{0}|X}^{\ast}(y_{1}\wedge y_{0}|X)$
$-W(F_{Y_{1}|X}^{\ast}(y_{1}\vee y_{0}|X),F_{Y_{0}|X}^{\ast}(y_{1}\wedge y_{0}|X))$
$-W(F_{Y_{1}|X}^{\ast}(y_{1}\wedge y_{0}|X),F_{Y_{0}|X}^{\ast}(y_{1}\vee y_{0}|X)),$
$A_{W}^{\ast}(S,X)$ $\equiv F_{Y_{1}|SX}^{\ast}(y_{1}\vee y_{0}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{1}\wedge y_{0}|S,X)$
$-W(F_{Y_{1}|SX}^{\ast}(y_{1}\vee y_{0}|S,X),F_{Y_{0}|SX}^{\ast}(y_{1}\wedge y_{0}|S,X))$
$-W(F_{Y_{1}|SX}^{\ast}(y_{1}\wedge y_{0}|S,X),F_{Y_{0}|SX}^{\ast}(y_{1}\vee y_{0}|S,X)).$
Taking expectations of (43) and (44) yields
$2\theta^{U}$ $=\mathbb{E}[\psi(Y_{1},Y_{1})]+\mathbb{E}[\psi(Y_{0},Y_{0})]-\int\int\mathbb{E}[A_{W}^{\ast}(X)]d\psi_{c}(y_{1},y_{0}),$
$2\theta_{U}$ $=\mathbb{E}[\psi(Y_{1},Y_{1})]+\mathbb{E}[\psi(Y_{0},Y_{0})]-\int\int\mathbb{E}[A_{W}^{\ast}(S,X)]d\psi_{c}(y_{1},y_{0}).$
Note that $\mathbb{E}[A_{W}^{\ast}(X)]$ and $\mathbb{E}[A_{W}^{\ast}(S,X)]$ can be expressed as
$\mathbb{E}[A_{W}^{\ast}(X)]$ $=F_{Y_{1}}^{\ast}(y_{1}\wedge y_{0})+F_{Y_{0}}^{\ast}(y_{1}\wedge y_{0})$
$-F^{\ast,(+)}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F^{\ast,(+)}(y_{1}\wedge y_{0},y_{1}\vee y_{0}),$
$\mathbb{E}[A_{W}^{\ast}(S,X)]$ $=F_{Y_{1}}^{\ast}(y_{1}\wedge y_{0})+F_{Y_{0}}^{\ast}(y_{1}\wedge y_{0})$
$-F_{(+)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F_{(+)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0}).$
Hence,
$\theta^{U}-\theta_{U}$ $=\frac{1}{2}\int\int(\mathbb{E}[A_{W}^{\ast}(S,X)]-\mathbb{E}[A_{W}^{\ast}(X)])d\psi_{c}(y_{1},y_{0})$
$=\frac{1}{2}\int\int(F^{\ast,(+)}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F_{(+)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0}))d\psi_{c}(y_{1},y_{0})$
$+\frac{1}{2}\int\int(F^{\ast,(+)}(y_{1}\wedge y_{0},y_{1}\vee y_{0})-F_{(+)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0}))d\psi_{c}(y_{1},y_{0}),$ (45)
where $\theta^{U}-\theta_{U}\geq 0$ holds by (42).
Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (42) and (45) that $\theta^{U}=\theta_{U}$ if and only if $F^{\ast,(+)}(y_{1},y_{0})=F_{(+)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Moreover, from equation (42), it follows that for $\psi_{c}$-almost every $(y_{1},y_{0})$, the condition $F^{\ast,(+)}(y_{1},y_{0})=F_{(+)}^{\ast}(y_{1},y_{0})$ holds if and only if
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\middle|X\right)\in\{0,1\}\text{ a.s.}$
Similarly, we can show that
$2\theta^{L}$ $=\mathbb{E}[\psi(Y_{1},Y_{1})]+\mathbb{E}[\psi(Y_{0},Y_{0})]-\int\int\mathbb{E}[A_{M}^{\ast}(X)]d\psi_{c}(y_{1},y_{0}),$
$2\theta_{L}$ $=\mathbb{E}[\psi(Y_{1},Y_{1})]+\mathbb{E}[\psi(Y_{0},Y_{0})]-\int\int\mathbb{E}[A_{M}^{\ast}(S,X)]d\psi_{c}(y_{1},y_{0}),$
where
$\mathbb{E}[A_{M}^{\ast}(X)]$ $=F_{Y_{1}}^{\ast}(y_{1}\wedge y_{0})+F_{Y_{0}}^{\ast}(y_{1}\wedge y_{0})$
$-F^{\ast,(-)}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F^{\ast,(-)}(y_{1}\wedge y_{0},y_{1}\vee y_{0}),$
$\mathbb{E}[A_{M}^{\ast}(S,X)]$ $=F_{Y_{1}}^{\ast}(y_{1}\wedge y_{0})+F_{Y_{0}}^{\ast}(y_{1}\wedge y_{0})$
$-F_{(-)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F_{(-)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0}).$
These results lead to
$\theta_{L}-\theta^{L}$ $=\frac{1}{2}\int\int(\mathbb{E}[A_{M}^{\ast}(X)]-\mathbb{E}[A_{M}^{\ast}(S,X)])d\psi_{c}(y_{1},y_{0})$
$=\frac{1}{2}\int\int(F_{(-)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F^{\ast,(-)}(y_{1}\vee y_{0},y_{1}\wedge y_{0}))d\psi_{c}(y_{1},y_{0})$
$+\frac{1}{2}\int\int(F_{(-)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0})-F^{\ast,(-)}(y_{1}\wedge y_{0},y_{1}\vee y_{0}))d\psi_{c}(y_{1},y_{0}),$ (46)
where $\theta_{L}-\theta^{L}\geq 0$ holds by (41).
Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (41) and (46) that $\theta^{L}=\theta_{L}$ if and only if $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, from equation (41), it follows that for $\psi_{c}$-almost every $(y_{1},y_{0})$, the condition $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ holds if and only if
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\middle|X\right)\in\{0,1\}\text{ a.s.}$
We have thus established the result under condition (a).
Under condition (b) of Proposition 4.1(i), it follows from equation (9) in Cambanis et al. (1976) that
$\theta^{L}(X)$ $=\mathbb{E}[\psi(Y_{1},\bar{y}_{0})|X]+\mathbb{E}[\psi(\bar{y}_{1},Y_{0})|X]-\psi(\bar{y}_{1},\bar{y}_{0})$
$+\int\int B_{M}^{\ast}(X)d\psi_{c}(y_{1},y_{0}),$ (47)
$\theta^{L}(S,X)$ $=\mathbb{E}[\psi(Y_{1},\bar{y}_{0})|S,X]+\mathbb{E}[\psi(\bar{y}_{1},Y_{0})|S,X]-\psi(\bar{y}_{1},\bar{y}_{0})$
$+\int\int B_{M}^{\ast}(S,X)d\psi_{c}(y_{1},y_{0}),$ (48)
where for all $(y_{1},y_{0})$,
$B_{M}^{\ast}(X)$ $\equiv M(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X))-\mathbf{1}\{\bar{y}_{1}<y_{1}\}F_{Y_{0}|X}^{\ast}(y_{0}|X)$
$-\mathbf{1}\{\bar{y}_{0}<y_{0}\}F_{Y_{1}|X}^{\ast}(y_{1}|X)+\mathbf{1}\{\bar{y}_{1}<y_{1}\}\mathbf{1}\{\bar{y}_{0}<y_{0}\},$
$B_{M}^{\ast}(S,X)$ $\equiv M(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X))-\mathbf{1}\{\bar{y}_{1}<y_{1}\}F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)$
$-\mathbf{1}\{\bar{y}_{0}<y_{0}\}F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+\mathbf{1}\{\bar{y}_{1}<y_{1}\}\mathbf{1}\{\bar{y}_{0}<y_{0}\}.$
Taking exceptions of (47) and (48) yields
$\theta^{L}$ $=\mathbb{E}[\psi(Y_{1},\bar{y}_{0})]+\mathbb{E}[\psi(\bar{y}_{1},Y_{0})]-\psi(\bar{y}_{1},\bar{y}_{0})+\int\int\mathbb{E}[B_{M}^{\ast}(X)]d\psi_{c}(y_{1},y_{0}),$
$\theta_{L}$ $=\mathbb{E}[\psi(Y_{1},\bar{y}_{0})]+\mathbb{E}[\psi(\bar{y}_{1},Y_{0})]-\psi(\bar{y}_{1},\bar{y}_{0})+\int\int\mathbb{E}[B_{M}^{\ast}(S,X)]d\psi_{c}(y_{1},y_{0}),$
where
$\mathbb{E}[B_{M}^{\ast}(X)]$ $=F^{\ast,(-)}(y_{1},y_{0})-\mathbf{1}\{\bar{y}_{1}<y_{1}\}F_{Y_{0}}^{\ast}(y_{0})$
$-\mathbf{1}\{\bar{y}_{0}<y_{0}\}F_{Y_{1}}^{\ast}(y_{1})+\mathbf{1}\{\bar{y}_{1}<y_{1}\}\mathbf{1}\{\bar{y}_{0}<y_{0}\}\mbox{\ \ and}$
$\mathbb{E}[B_{M}^{\ast}(S,X)]$ $=F_{(-)}^{\ast}(y_{1},y_{0})-\mathbf{1}\{\bar{y}_{1}<y_{1}\}F_{Y_{0}}^{\ast}(y_{0})$
$-\mathbf{1}\{\bar{y}_{0}<y_{0}\}F_{Y_{1}}^{\ast}(y_{1})+\mathbf{1}\{\bar{y}_{1}<y_{1}\}\mathbf{1}\{\bar{y}_{0}<y_{0}\},$
for all $(y_{1},y_{0})$.
Then it follows that
$\theta_{L}-\theta^{L}$ $=\int\int(\mathbb{E}[B_{M}^{\ast}(S,X)]-\mathbb{E}[B_{M}^{\ast}(X)])d\psi_{c}(y_{1},y_{0})$
$=\int\int(F_{(-)}^{\ast}(y_{1},y_{0})-F^{\ast,(-)}(y_{1},y_{0}))d\psi_{c}(y_{1},y_{0}),$ (49)
where $\theta_{L}-\theta^{L}\geq 0$ holds from (41).
Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (41) and (49) that $\theta^{L}=\theta_{L}$ if and only if $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, for $\psi_{c}$-almost every $(y_{1},y_{0})$, the condition $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ holds if and only if
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\middle|X\right)\in\{0,1\}\text{ a.s.}$
For the upper bounds under condition (b), by a similar argument, we can show that $\theta^{U}=\theta_{U}$ if and only if $\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\middle|X\right)\in\{0,1\}$ a.s. Thus, the result under condition (b) has now been established. ∎
Proof of Proposition 4.3.
Fix $\delta$. We first prove part (i). Under the stated conditions and by the definition of $\widetilde{\Theta}_{I}$, Theorem 3.5(i) in Fan et al. (2017) implies $\widetilde{\Theta}_{I}=[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$. Proposition 3.1 then yields $\Theta_{I}=\widetilde{\Theta}_{I}=[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$.
For part (ii), under the stated conditions and by the definition of $\widetilde{\Theta}_{IC}$, applying Theorem 3.5(i) in Fan et al. (2017) with $X$ replaced by $(S,X)$ gives $\widetilde{\Theta}_{IC}=[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)]$. Theorem 3.3 then yields $\Theta_{IC}=\widetilde{\Theta}_{IC}=[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)]$. ∎
Proof of Theorem 4.4.
We provide a proof for the lower bounds. The proof for the upper bounds is analogous and therefore omitted. By the definitions of $F_{\varphi}^{L}(\delta)$ and $F_{L,\varphi}(\delta)$ and by Jensen’s inequality, we have
$F_{\varphi}^{L}(\delta)$ $=\mathbb{E}\left[\sup_{y\in\mathcal{Y}_{1}}\max\left\{F_{Y_{1}|X}^{\ast}(y|X)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|X)-1,0\right\}\right]$
$=\mathbb{E}\left[\sup_{y\in\mathcal{Y}_{1}}\max\left\{\mathbb{E}\bigl[F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1|X\bigr],0\right\}\right]$
$\leq\mathbb{E}\left[\sup_{y\in\mathcal{Y}_{1}}\max\left\{F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1,0\right\}\right]$
$=F_{L,\varphi}(\delta).$ (50)
Note that $Y_{d}$ ($d=0,1$) are assumed to be continuous random variables. Hence $F_{Y_{d}|SX}^{\ast}(y|S,X)$ and $F_{Y_{d}|X}^{\ast}(y|X)$ are continuous, and thus $\sup_{y\in\mathcal{Y}_{1}}F_{Y_{1}|SX}^{\ast}(y|S,X)=1$ and $\sup_{y\in\mathcal{Y}_{1}}F_{Y_{1}|X}^{\ast}(y|X)=1$. This implies that
$\sup_{y\in\mathcal{Y}_{1}}\{F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1\}\geq 0\mbox{ a.s.},$
and
$\sup_{y\in\mathcal{Y}_{1}}\{F_{Y_{1}|X}^{\ast}(y|X)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|X)-1\}\geq 0\mbox{ a.s.}$
Therefore, equation (50) simplifies to
$F_{\varphi}^{L}(\delta)$ $=\mathbb{E}\left[\sup_{y\in\mathcal{Y}_{1}}\left\{F_{Y_{1}|X}^{\ast}(y|X)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|X)-1\right\}\right]$
$=\mathbb{E}\left[\sup_{y\in\mathcal{Y}_{1}}\left\{\mathbb{E}\bigl[F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1|X\bigr]\right\}\right]$
$\leq\mathbb{E}\left[\sup_{y\in\mathcal{Y}_{1}}\left\{F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1\right\}\right]$
$=F_{L,\varphi}(\delta).$ (51)
Let $G_{\varphi}(y,s,x)=F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1$. Then $\sup_{y\in\mathcal{Y}_{1}}\mathbb{E}[G(y,S,X)|X=x]=\mathbb{E}[G(\bar{y}(x),S,X)|X=x]$ and it follows from (51) that $F_{\varphi}^{L}(\delta)=F_{L,\varphi}(\delta)$ if and only if $\mathbb{E}[\sup_{y\in\mathcal{Y}_{1}}G(y,S,X)|X=x]=\mathbb{E}[G(\bar{y}(x),S,X)|X=x]$ for almost all $x\in\mathcal{X}$.
Since $\sup_{y\in\mathcal{Y}_{1}}G(y,s,x)\geq G(\bar{y}(x),s,x)$ for all $(s,x)\in\{0,1\}\times\mathcal{X}$, it follows that
$\mathbb{E}[\sup_{y\in\mathcal{Y}_{1}}G(y,S,X)|X=x]=\mathbb{E}[G(\bar{y}(x),S,X)|X=x]$
for almost all $x\in\mathcal{X}$ if and only if $\sup_{y\in\mathcal{Y}_{1}}G(y,s,x)=G(\bar{y}(x),s,x)$ for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$; that is, $G_{\varphi}(y,0,x)$ and $G_{\varphi}(y,1,x)$ attain their maxima at the common point $\bar{y}(x)$ for almost all $x\in\mathcal{X}$. ∎
Proof of Proposition 5.1.
Let $\mathcal{F}$ denote the class of all distribution functions of the variables $(Y_{1},Y_{0},Y,D,S,G,X)$. For any generic subclass $\widetilde{\mathcal{F}}\subseteq\mathcal{F}$, let $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}=\{F_{Y_{1}Y_{0}SX}:F\in\widetilde{\mathcal{F}}\}$ be the corresponding set of marginal distributions of $(Y_{1},Y_{0},S,X)$. Next, let $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ denote the set of distribution functions of $(Y_{1},Y_{0},S,X)$ that satisfy constraints (16)–(18):
[…]
Define
$\widetilde{\Theta}_{IC}^{\dagger}\equiv\bigl\{\mathbb{E}_{F_{Y_{1}Y_{0}SX}}[\psi(Y_{1},Y_{0})]:F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}\bigr\}.$
We also introduce the functional $\Gamma:\mathcal{F}_{Y_{1}Y_{0}SX}\rightarrow\mathbb{R}$ defined by $\Gamma(F_{Y_{1}Y_{0}SX})\equiv\mathbb{E}_{F_{Y_{1}Y_{0}SX}}[\psi(Y_{1},Y_{0})]$. It then follows that $\widetilde{\Theta}_{IC}^{\dagger}=\bigl\{\Gamma(F_{Y_{1}Y_{0}SX}):F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}\bigr\}.$
Since all restrictions (16)–(18) are linear and $\mathcal{F}_{Y_{1}Y_{0}SX}$ is convex, $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ is convex as well. Moreover, because $\Gamma:\mathcal{F}_{Y_{1}Y_{0}SX}\rightarrow\mathbb{R}$ is linear, the convexity of $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ implies that $\widetilde{\Theta}_{IC}^{\dagger}$ is convex. Hence its closure is
$\left[\inf_{F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}}\Gamma(F_{Y_{1}Y_{0}SX}),\sup_{F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}}\Gamma(F_{Y_{1}Y_{0}SX})\right]=[\theta_{L}^{\ast},\theta_{U}^{\ast}].$
We next show that $\widetilde{\Theta}_{IC}^{\dagger}=\Theta_{IC}^{\dagger}$. For conditional joint distributions $F_{Y_{1}Y_{0}|SX}$, we consider the following conditions for all $(s,x)$:
$F_{Y_{d}|Y_{d^{\prime}}\leq y,S=s,X=x}(t)\geq F_{Y_{d}|Y_{d^{\prime}}\leq y^{\prime},S=s,X=x}(t)$
[…]
and
$\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=1,X=x)\geq\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=0,X=x)\mbox{ for all }c\in\mathbb{R}$ (53)
Conditions (52) and (53) are equivalent to Assumptions 5.1 and 5.2, respectively, conditional on $(S,X)=(s,x)$.
For each $x\in\mathcal{X}$, we define the following class of pairs of conditional copula functions:
$\widetilde{\mathcal{C}}_{x}\equiv\bigg\{(C(\cdot,\cdot|0,x),C(\cdot,\cdot|1,x))\in\mathcal{C}^{2}:\;F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|s,x):=C(F_{Y_{1}|SX}^{\ast}(\cdot|s,x),F_{Y_{0}|SX}^{\ast}(\cdot|s,x))$
[…]
That is, $\widetilde{\mathcal{C}}_{x}$ consists of all pairs of conditional copula functions that generate conditional joint distributions $F_{Y_{1}Y_{0}|SX}$ satisfying conditions (52) and (53). Note that $\widetilde{\mathcal{C}}_{x}$ is nonempty, since there exists a pair $(C^{\ast}(\cdot,\cdot|0,x),C^{\ast}(\cdot,\cdot|1,x))\in\widetilde{\mathcal{C}}_{x}$ such that
$F_{Y_{1}Y_{0}|SX}^{\ast}(\cdot,\cdot|s,x)=C^{\ast}(F_{Y_{1}|SX}^{\ast}(\cdot|s,x),F_{Y_{0}|SX}^{\ast}(\cdot|s,x)),$
for all $(s,x)$.
Then $\widetilde{\Theta}_{IC}^{\dagger}$ can be expressed as
$\widetilde{\Theta}_{IC}^{\dagger}$ $=\left\{\theta:\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)|S,X\bigr)\right]\right$.
[…]
We first show that $\Theta_{IC}^{\dagger}\subseteq\widetilde{\Theta}_{IC}^{\dagger}$. Fix any $\theta\in\Theta_{IC}^{\dagger}$. By the definition of $\Theta_{IC}^{\dagger}$, there exists a distribution function $F\in\widetilde{\mathcal{F}}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Since $F_{YDGX}=F_{YDGX}^{\ast}$ and $F$ satisfies Assumptions 2.12.5, Lemma 3.2 implies that $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$ for $d=0,1$.
By Sklar’s theorem and the definition of $\widetilde{\mathcal{C}}_{x}$, the equalities $F_{Y_{d}\mid SX}=F_{Y_{d}\mid SX}^{\ast}$ for $d=0,1$ guarantee the existence of a pair of conditional copula functions $(C(\cdot,\cdot\mid 0,x),C(\cdot,\cdot\mid 1,x))\in\widetilde{\mathcal{C}}_{x}$ such that
$F_{Y_{1}Y_{0}|SX}(\cdot,|s,x)=C(F_{Y_{1}|SX}(\cdot|s,x),F_{Y_{0}|SX}(\cdot|s,x)|s,x),$
for almost all $(s,x)$.
It then follows that
$\theta$ $=\mathbb{E}_{F_{Y_{1}Y_{0}}}\left[\psi(Y_{1},Y_{0})\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|SX}(y_{1},y_{0}|S,X)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}(y_{1}|S,X),F_{Y_{0}|SX}(y_{0}|S,X)\big|S,X\bigr)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right]$
$\in\widetilde{\Theta}_{IC}^{\dagger},$
where the fourth line uses $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, and the last line follows from (54). Thus, we have $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$. Since this argument holds for any $\theta\in\Theta_{IC}^{\dagger}$, it follows that $\Theta_{IC}^{\dagger}\subseteq\widetilde{\Theta}_{IC}^{\dagger}$.
We next show that $\widetilde{\Theta}_{IC}^{\dagger}\subseteq\Theta_{IC}^{\dagger}$. Fix any $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$. By the definition of $\widetilde{\Theta}_{IC}^{\dagger}$, there exists a pair of conditional copula functions $(C(\cdot,\cdot|s,x),C(\cdot,\cdot|s,x))\in\widetilde{\mathcal{C}}_{x}$ such that
$\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right].$
We now show that there exists a distribution function $F\in\widetilde{\mathcal{F}}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ hierarchically, defined by
$F_{DGX}=F_{DGX}^{\ast};$ (55)
$F_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|D,\mathrm{obs},X)=C\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|D,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|D,\mathrm{obs},X)\middle|D,X\right);$ (56)
$F_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|GX}(y_{1},y_{0}|\mathrm{obs},X);;$ (57)
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{obs},X)=\mathbf{1}\{D\leq s\};$ (58)
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{exp},X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X);$ (59)
$F_{Y|Y_{1}Y_{0}DSGX}(y|Y_{1},Y_{0},D,S,G,X)=\mathbf{1}\{DY_{1}+(1-D)Y_{0}\leq y\}.$ (60)
We will show that $F\in\widetilde{\mathcal{F}}^{\ast}$ and that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 and 5.15.2 (with $F^{\ast}$ replaced by $F$) and that (ii) $F_{YDGX}=F_{YDGX}^{\ast}$.
We begin by verifying condition (i). By the same argument as in the proof of Theorem 3.3, it follows that that $F$ satisfies Assumptions 2.12.5, or equivalently $F\in\mathcal{F}^{\ast}$. Applying Lemma C.1(i) to (56), we obtain
$F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|S,X)=C\left(F_{Y_{1}|SX}^{\ast}(\cdot|S,X),F_{Y_{0}|SX}^{\ast}(\cdot|S,X)\middle|S,X\right)\mbox{ a.s}.$
By the definition of $\widetilde{\mathcal{C}}_{X}$, $F_{Y_{1}Y_{0}|SX}$ satisfies conditions (52) and (53). Hence, $F$ also satisfies Assumptions 5.1 and 5.2. In summary, we have shown that $F$ satisfies condition (i).
Condition (ii), namely $F_{YDGX}=F_{YDGX}^{\ast}$, follows by the same argument as in the proof of Theorem 3.3.
We subsequently show that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. It follows that
$\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ $=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|SX}(y_{1},y_{0}|S,X)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dF_{Y_{1}Y_{0}|DGX}(y_{1},y_{0}|S,\mathrm{obs},X)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|S,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|S,\mathrm{obs},X)\middle|S,X\right)\right]$
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right]$
$=\theta,$
where the second and fourth equalities follow from Lemma C.1(i), the third from (56), and the last from (32). The fifth equality follows because (55) together with Lemma C.1(ii) implies $F_{SX}=F_{SX}^{\ast}$. Thus, we have established that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$.
Consequently, we have shown that $F\in\widetilde{\mathcal{F}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$; thus, $\theta\in\Theta_{IC}^{\dagger}$. Since this argument holds for any $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$, we conclude that $\widetilde{\Theta}_{IC}^{\dagger}\subseteq\Theta_{IC}^{\dagger}$. ∎